### v4

# Definability of mad families of vector spaces and two local Ramsey theories

Abstract.
Let $E$ be a vector space over a countable field of dimension $\aleph_{0}$. Two infinite-dimensional subspaces $V,W\subseteq E$ are almost disjoint if $V\cap W$ is finite-dimensional. This paper provides some improvements on results about the definability of maximal almost disjoint families (mad families) of subspaces in [S19]. We construct a full mad family of block subspaces in $\mathsf{ZFC}$, answering a problem by Smythe in the positive. A variant of this construction shows that there exists a completely separable mad family of block subspaces in $\mathsf{ZFC}$. We also discuss the abstract Mathias forcing introduced by Di Prisco-Mijares-Nieto in [DMN15], and apply it to show that in the Solovay’s model obtained by the collapse of a Mahlo cardinal, there are no full mad families of block subspaces over $\mathbb{F}_{2}$.
1. Introduction
Let $E$ be a vector space over a countable field $\mathbb{F}$ of dimension $\aleph_{0}$. In [S19], Smythe studied almost disjoint families infinite-dimensional subspaces $V,W\subseteq E$, i.e. families of infinite-dimensional subspaces of $E$ which pairwise intersections are finite-dimensional. Our paper focuses on the following problems raised by Smythe in [S19].
Problem 1.1.
Can mad families of subspaces be definable? In particular:
  1. (1)
    Are there analytic mad families of subspaces?
  2. (2)
    Assuming the existence of large cardinals, are there mad families of subspaces in $\mathbf{L}(\mathbb{R})$?
Horowitz-Shelah proved in [HS22] that there are no analytic mad families of subspaces over $\mathbb{F}_{2}$, but the first problem remains largely open for other fields. Smythe remarked in [S19] that if there is a supercompact cardinal, then every set in Solovay’s model is $\mathcal{H}$-strategically Ramsey, so there are no full mad families (the definition of a full mad family will be introduced in the subsequent paragraph). Strategically Ramsey subsets of $E^{[\infty]}$ were introduced in [R08] and [R09] to study the combinatorics of Banach space theory and dichotomies, and Smythe introduced a localised version of the strategically Ramsey property in [S18]. To the author’s knowledge, the precise large cardinal strength needed for the non-existence of mad families in $\mathbf{L}(\mathbb{R})$ is mostly unknown.
Inspired by Mathias’ original proof of the non-existence of analytic mad families of $\omega]^{\omega}$, Smythe approached this problem using a local Ramsey theoretic approach in [S19], replacing $\mathcal{H}$-Ramsey with $\mathcal{H}$-strategically Ramsey. Given an almost disjoint family $\mathcal{A}$ of vector spaces, he extended Mathias’ definition of the “happy family” $\mathfrak{F}_{\mathcal{A}}$ of subsets of $\omega$, to a “family” $\mathcal{H}(\mathcal{A})$ of subspaces of $E$. His approach requires $\mathcal{H}(\mathcal{A})$ to be full, a property which closely resembles the pigeonhole property of coideals of $\omega]^{\omega}$. Calling a mad family $\mathcal{A}$ full if $\mathcal{H}(\mathcal{A})$ is full, Smythe proved the following theorem:
Theorem 1.2 (Smythe [S19]).
There exist no analytic full mad families of block subspaces.
Block subspaces are subspaces with a basis of vectors with strictly increasing support (see §2 for a precise definition). Block subspaces are more combinatorially tame, and since every infinite-dimensional subspace contains an infinite-dimensional block subspace, an almost disjoint family of subspaces is maximal with respect to block subspaces iff it is maximal.
While it is clear from Smythe’s proof that the fullness property plays an important role in understanding the definability of mad families of subspaces, the extent which mad families of subspaces satisfy such property remains poorly understood. In particular, Smythe asked the following question in [S19]:
Problem 1.3 (Smythe, [S19]).
Does $\mathsf{ZFC}$ prove the existence of a full mad family of subspaces? Even more, does $\mathsf{ZFC}$ prove that every mad family of subspaces is full?
In the same paper, Smythe proved that assuming $\mathfrak{p}=\mathfrak{c}$, there exists a full mad family of block subspaces.
Another obstacle to understanding the definability of mad families of subspaces is the restriction to mad families of block subspaces. It is unknown if Theorem 1.2 holds for arbitrary mad families of subspaces (except for when $|\mathbb{F}|=2$). Furthermore, consider the following two cardinals:
[…]
It is unknown if they are equal in $\mathsf{ZFC}$.
Problem 1.4 (Smythe, [S19]).
Do we lose any generality when restricting mad families to block subspaces? In particular, is $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}=\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ true in $\mathsf{ZFC}$?
The first result in the paper answers the first part of Problem 1.1 in the positive.
Theorem 1.5.
There exists a full mad family of block subspaces.
Our method is inspired by a variant of the construction of a completely separable mad family of $\omega]^{\omega}$ in [MRS14] assuming $\mathfrak{s}\leq\mathfrak{a}$, provided by Guzmán in [G21]. Construction of a completely separable mad family of $\omega]^{\omega}$ has been extensively studied in the past two decades. Shelah attained a breakthrough in [Sh11], where he showed that a completely separable mad family assuming either $\mathfrak{c}<\aleph_{\omega}$, or there is no inner model with a measurable cardinal. However, the existence of a completely separable mad family in $\mathsf{ZFC}$ alone remains open. We shall show that the natural generalisation of a completely separable mad family to vector spaces (see Definition 5.6) exists in $\mathsf{ZFC}$.
Theorem 1.6.
There exists a completely separable mad family of block subspaces.
Our constructions for Theorem 1.5 and 1.6 utilises various results from the local Ramsey theory of block subspaces developed by Smythe in [S18] and [S19], along with a few additional results that we shall prove.
The second part in this paper addresses the existence of full mad families of block subspaces over $\mathbb{F}_{2}$ in Solovay’s model. We utilise the abstract Mathias forcing developed by Di Prisco-Mijares-Nieto in [DMN15] - the abstract Mathias forcing was developed to expand the local Ramsey theory of $\omega]^{\omega}$ to topological Ramsey spaces (i.e. closed triples $(\mathcal{R},\leq,r)$ satisfying the axioms A1-A4 introduced by Todorčević in [T10]). For example, if $E$ is a countable vector space over $\mathbb{F}_{2}$, then the pigeonhole principle holds, so it is a topological Ramsey space.
Let $\operatorname{Coll}(\omega,<\!\kappa)$ be the Lévy collapse, which collapses the uncountable cardinal $\kappa$ to $\omega_{1}$. Using the abstract Mathias forcing, we may apply the abstract Mathias forcing to obtain the following variant of a celebrated theorem of Mathias:
Theorem 1.7.
Let $\kappa$ be a Mahlo cardinal, and let $G$ be a $\operatorname{Coll}(\omega,<\!\kappa)$-generic filter. Suppose that $E$ is a countable vector space over $\mathbb{F}_{2}$. If $\mathcal{H}\in\mathbf{V}[G]$ is a selective coideal on $E^{[\infty]}$, and $\mathcal{X}\in\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$ is a subset of $E^{[\infty]}$, then $\mathcal{X}$ is $\mathcal{H}$-Ramsey.
As a consequence, we obtain the following corollary:
Corollary 1.8.
Let $\kappa$ be a Mahlo cardinal, and let $G$ be a $\operatorname{Coll}(\omega,<\!\kappa)$-generic filter. Then there are no full mad families of block subspaces over $\mathbb{F}_{2}$ in $\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$.
Our paper is organised as follows:
  1. 2)
    We provide a concise introduction to various definitions and notations to block subspaces of $E$.
  2. 3)
    We review the local Ramsey theory of block subspaces developed by Smythe, i.e. combinatorial properties of semicoideals of block subspaces (which Smythe calls families). We introduce the notion of a selective semicoideal of block subspaces, a strengthening of the (p)-property.
  3. 4)
    We introduce some definitions and basic results about almost disjoint families of (block) subspaces. We show that $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ is uncountable, addressing a remark by Smythe in [S19]. We also show that the semicoideal $\mathcal{H}^{-}(\mathcal{A})$ is selective for any almost disjoint family $\mathcal{A}$ - the non-requirement for $\mathcal{A}$ to be maximal is crucial in the proof of Theorem 1.5 and 1.6.
  4. 5)
    We apply various lemmas and results proven in §3 and §4 to provide a detailed proof of Theorem 1.5 and 1.6.
  5. 6)
    We provide an overview of the axioms of topological Ramsey spaces A1-A4 introduced by Todorčević, and the abstract local Ramsey theory developed in [DMN15]. We show that given a closed triple $(\mathcal{R},\leq,r)$ satisfying A1-A4, if $\mathcal{H}\subseteq\mathcal{R}$ is semiselective then $\mathbb{M}_{\mathcal{H}}$ satisfies the Mathias property and the Prikry property. This was proven in [DMN15], but they assumed that the almost reduction relation $\leq^{*}$ is transitive (which is not generally true, such as for the topological Ramsey space of strong subtrees). We provide a proof of this result without this assumption.
  6. 7)
    We study the intersection between the local Ramsey theory of block subspaces and abstract local Ramsey theory when considering block subspaces over $\mathbb{F}_{2}$. We show that various notions and results introduced in §3 and §6 coincide in the setting of block subspaces over $\mathbb{F}_{2}$ and provide a proof of Theorem 1.7 and Corollary 1.8.
  7. 8)
    We conclude the paper with several further remarks and open problems.
2. Vector spaces and block subspaces
We introduce some basic definitions and properties of vector spaces and block subspaces.
2.1. Setup and notations
Let $E$ be a vector space over a countable field $\mathbb{F}$, with a countable Hamel basis $(e_{n})_{n<\omega}$. For each $x\in E$, we may write $x:=\sum_{n<\omega}\lambda_{n}(x)e_{n}$, where $\lambda_{n}\in\mathbb{F}$. This allows us to define the support of a vector $x\in E$ by:
$\operatorname{supp}(x):=\{n<\omega:\lambda_{n}(x)\neq 0\}.$
This, in turn, gives us a partial order << on $E$, where for any two non-zero vectors $x,y\in E$:
$x<y\iff\max(\operatorname{supp}(x))<\min(\operatorname{supp}(y)).$
A block sequence of vectors in $E$ is thus a <<-increasing sequence. Note that any block sequence of vectors is linearly independent.
Definition 2.1.
An infinite-dimensional subspace $V\subseteq E$ is a block subspace if $V$ has a block basis, i.e. an infinite block sequence $(x_{n})_{n<\omega}$ such that $V=\operatorname{span}\{x_{n}:n<\omega\}$.
Since such a block basis is unique up to scaling, we may conflate block subspaces of $E$ with infinite block sequences.
Notation 2.2.
  • We let $E^{[\infty]}$ (resp. $E^{[<\infty]}$) denote the set of infinite (resp. finite) block sequences of vectors in $E$.
  • $E^{[\leq\infty]}:=E^{[\infty]}\cup E^{[<\infty]}$.
  • If $A=(x_{n})_{n<\omega}\in E^{[\infty]}$ and $N<\omega$, we write $A/N:=(x_{n})_{n\geq N}$.
  • If $A=(x_{n})_{n<\omega}\in E^{[\infty]}$ and $n<\omega$, we write $r_{n}(A):=(x_{i})_{i<n}$.
  • If $A\in E^{[<\infty]}$ and $a\in E^{[<\infty]}$ we write $d_{A}(a):=\min\{N<\omega:\max(\operatorname{supp}(a))<A/N\}$.
  • If $A\in E^{[\infty]}$ and $a\in E^{[<\infty]}$, then $A/a:=A/d_{A}(a)$.
  • If $(x_{n})_{n<\alpha}\in E^{[\leq\infty]}$, we write:
    $\langle(x_{n})_{n<\alpha}\rangle:=\operatorname{span}\{x_{n}:n<\alpha\}.$
  • If $A,B\in E^{[\infty]}$, we write $A\leq B$ iff $\langle A\rangle\subseteq\langle B\rangle$.
  • If $A,B\in E^{[\infty]}$, we write $A\leq^{*}B$ iff $A/N\leq B$ for some $N<\omega$.
Block sequences/subspaces were introduced as they are more combinatorially well-behaved than arbitrary infinite-dimensional subspaces. The focus on block sequences is further justified by the observation that every infinite-dimensional subspace of $E$ contains a block subspace. This will be apparent once we have introduced the relevant definitions and notations for almost disjoint families of subspaces.
2.2. Intersection of block subspaces
Given $A,B\in E^{[\infty]}$, if $\langle A\rangle\cap\langle B\rangle$ is infinite-dimensional, then by the remark above it contains a block subspace. The following lemma shows that $\langle A\rangle\cap\langle B\rangle$ is itself a block subspace. This observation is central in ensuring that the proof of Theorem 1.5 and 1.6 do not require additional set-theoretic hypotheses.
Lemma 2.3.
Let $A,B\in E^{[\infty]}$ be such that $\langle A\rangle\cap\langle B\rangle$ is infinite-dimensional. Then there exists some $C\in E^{[\infty]}$ such that $\langle C\rangle=\langle A\rangle\cap\langle B\rangle$.
Proof.
We write $A=(x_{n})_{n<\omega}$ and $B=(y_{n})_{n<\omega}$. For each $v\in\langle A\rangle\cap\langle B\rangle$, we let:
$\operatorname{supp}_{A}(v)$ $:=\{n<\omega:\operatorname{supp}(x_{n})\subseteq\operatorname{supp}(v)\},$
$\operatorname{supp}_{B}(v)$ $:=\{n<\omega:\operatorname{supp}(y_{n})\subseteq\operatorname{supp}(v)\}.$
We remark that for all $v,w\in\langle A\rangle\cap\langle B\rangle$:
$\operatorname{supp}(v)\subseteq\operatorname{supp}(w)$ $\iff\operatorname{supp}_{A}(v)\subseteq\operatorname{supp}_{A}(w)$
$\iff\operatorname{supp}_{B}(v)\subseteq\operatorname{supp}_{B}(w).$
For an arbitrary $v\in E$ and $k\in\operatorname{supp}(v)$, we also write $v(k):=e_{k}^{*}(v)\in\mathbb{F}$ to be the unique coefficient such that $v=u+v(k)e_{k}+w$ for some vectors $u<e_{k}<w$ (or $u=0$ or $w=0$).
Claim 1.
Let $v_{0},v_{1}\in\langle A\rangle\cap\langle B\rangle$, and suppose that $\operatorname{supp}_{A}(v_{0})\cap\operatorname{supp}_{A}(v_{1})\neq\emptyset$. Then there exists some $w\in\langle A\rangle\cap\langle B\rangle$ such that $\operatorname{supp}_{A}(w)=\operatorname{supp}_{A}(v_{0})\cap\operatorname{supp}_{A}(v_{1})$.
Proof.
Let $v_{0},v_{1}\in\langle A\rangle\cap\langle B\rangle$ be such that $\operatorname{supp}_{A}(v_{0})\cap\operatorname{supp}_{A}(v_{1})\neq\emptyset$. We may write:
$v_{0}$ $=\sum_{i\in\operatorname{supp}_{A}(v_{0})}\lambda_{i}^{0}x_{i}=\sum_{j\in\operatorname{supp}_{B}(v_{0})}\mu_{j}^{0}y_{j},$ (2.1)
$v_{1}$ $=\sum_{i\in\operatorname{supp}_{A}(v_{1})}\lambda_{i}^{1}x_{i}=\sum_{j\in\operatorname{supp}_{B}(v_{1})}\mu_{j}^{1}y_{j}.$ (2.2)
where $\lambda_{i}^{0},\lambda_{i}^{1},\lambda_{j}^{0},\lambda_{j}^{1}$ are all non-zero. We let:
$w:=\sum_{i\in\operatorname{supp}_{A}(v_{0})\cap\operatorname{supp}_{A}(v_{1})}\lambda_{i}^{0}x_{i},$
$w^{\prime}:=\sum_{j\in\operatorname{supp}_{B}(v_{0})\cap\operatorname{supp}_{B}(v_{1})}\mu_{j}^{0}y_{j}.$
It’s clear that $\operatorname{supp}_{A}(w)=\operatorname{supp}_{A}(v_{0})\cap\operatorname{supp}_{A}(v_{1})$. We shall show that $w=w^{\prime}$. This implies that that $w\in\langle A\rangle\cap\langle B\rangle$, proving the claim.
By the symmetry of the argument, it suffices to show that $\operatorname{supp}(w)\subseteq\operatorname{supp}(w^{\prime})$, and $w(k)=w^{\prime}(k)$ for all $k\in\operatorname{supp}(w)$. Let $k\in\operatorname{supp}(w)$, so $k\in\operatorname{supp}(x_{i})$ for some $i\in\operatorname{supp}_{A}(v_{0})\cap\operatorname{supp}_{A}(v_{1})$. By the second inequalities in (2.1) and (2.2), we have that $k\in\operatorname{supp}(y_{j})$ for some $j\in\operatorname{supp}_{B}(v_{0})$, and $k\in\operatorname{supp}(y_{j^{\prime}})$ for some $j^{\prime}\in\operatorname{supp}_{B}(v_{1})$. Since $B$ is a block sequence, $k\in\operatorname{supp}(y_{j})\cap\operatorname{supp}(y_{j^{\prime}})$ implies that $j=j^{\prime}$. Therefore, $j\in\operatorname{supp}_{B}(v_{0})\cap\operatorname{supp}_{B}(v_{1})$, so $k\in\operatorname{supp}(w^{\prime})$. By (2.1), we can conclude that:
$w(k)=\lambda_{i}^{0}x_{i}(k)=v_{0}(k)=\mu_{j}^{0}y_{j}(k)=w^{\prime}(k),$
which completes the proof. ∎
Let $K\subseteq\omega$ be the set of all $n$ such that $n\in\operatorname{supp}_{A}(v)$ for some $v\in\langle A\rangle\cap\langle B\rangle$. Since $\langle A\rangle\cap\langle B\rangle$ is infinite-dimensional, $K$ is infinite. Define a relation $\sim$ on $K$ such that for all $m,n\in K$:
$m\sim n\iff\forall v\in\langle A\rangle\cap\langle B\rangle(m\in\operatorname{supp}_{A}(v)\leftrightarrow n\in\operatorname{supp}_{A}(v)).$
It’s clear from the definition that $\sim$ is an equivalence relation, and that every equivalence class is finite.
Claim 2.
For each $n\in K$, there exists some $v\in\langle A\rangle\cap\langle B\rangle$ such that $\operatorname{supp}_{A}(v)=[n]_{\sim}$.
Proof.
It follows from the definition of $\sim$ that:
[…]
However, since $\operatorname{supp}_{A}(u)$ is finite for each $u\in\langle A\rangle\cap\langle B\rangle$, we may find $u_{0},\dots,u_{m-1}\in\langle A\rangle\cap\langle B\rangle$ such that $n]_{\sim}=\bigcap_{i<m}\operatorname{supp}_{A}(u_{i})$. Then the existence of such a $v$ follows from Claim 1. ∎
Thus, for each $n$ we let $z_{n}\in\langle A\rangle\cap\langle B\rangle$ be a fixed vector such that $\operatorname{supp}_{A}(z_{n})=[n]_{\sim}$.
Claim 3.
Let $m,n,l\in K$ be such that $m<n<l$, and suppose that $m\sim l$. Then $m\sim n\sim l$.
Proof.
Suppose otherwise, so there exist $m<n<l$ in $K$ such that $m\sim l$ but $n\notin[m]_{\sim}$. We write:
$z_{m}$ $=\sum_{i\in\operatorname{supp}_{A}(z_{m})}\lambda_{i}x_{i}=\sum_{j\in\operatorname{supp}_{B}(z_{m})}\mu_{j}y_{j},$
where $\lambda_{i},\mu_{j}$ are non-zero. Let $i_{0}:=\max\{i\in\operatorname{supp}_{A}(z_{m}):i<n\}$ and $i_{1}:=\min\{i\in\operatorname{supp}_{A}(z_{m}):i>n\}$, and let $k_{0}:=\max(\operatorname{supp}(x_{i_{0}}))$ and $k_{1}:=\min(\operatorname{supp}(x_{i_{1}}))$. We define:
$w:=\sum_{\begin{subarray}{c}i\in\operatorname{supp}_{A}(z_{m})\\ \max(\operatorname{supp}(x_{i}))\leq k_{0}\end{subarray}}\lambda_{i}^{m}x_{i}.$
We shall show that $w\in\langle A\rangle\cap\langle B\rangle$. This gives us the required contradiction, as $m\in\operatorname{supp}_{A}(w)$ but $l\not\in\operatorname{supp}_{A}(w)$.
Observe that for each $j\in\operatorname{supp}_{B}(z_{m})$, either $k_{0}<\min(\operatorname{supp}(y_{j}))$ or $\max(\operatorname{supp}(y_{j}))<k_{1}$ – otherwise, if $j\in\operatorname{supp}_{B}(z_{m})$ is such that $\min(\operatorname{supp}(y_{j}))\leq k_{0}<k_{1}\leq\max(\operatorname{supp}(y_{j}))$, then by the block sequence property of $B$, we have that for all $j^{\prime}\neq j$, either $\max(\operatorname{supp}(y_{j^{\prime}}))<k_{0}$ or $k_{1}<\min(\operatorname{supp}(y_{j^{\prime}}))$. Since $k_{0}<\min(\operatorname{supp}(x_{n}))\leq\max(\operatorname{supp}(x_{n}))<k_{1}$ and $j\notin\operatorname{supp}_{B}(z_{n})$, this implies that:
$\operatorname{supp}(x_{n})\cap\bigcup_{j\in\operatorname{supp}_{B}(z_{n})}\operatorname{supp}(y_{j})=\emptyset,$
which contradicts that $\operatorname{supp}(x_{n})\subseteq\operatorname{supp}(z_{n})$.
Note that for all $k\in\operatorname{supp}(z_{m})$, either $k\leq k_{0}$ or $k_{1}\leq k$. Thus, for all $j\in\operatorname{supp}_{B}(z_{m})$, since $\operatorname{supp}(y_{j})\subseteq\operatorname{supp}(z_{m})$, if $\max(\operatorname{supp}(y_{j}))<k_{1}$ then $\max(\operatorname{supp}(y_{j}))\leq k_{0}$. Similarly, if $k_{0}<\min(\operatorname{supp}(y_{j}))$ then $k_{1}\leq\min(\operatorname{supp}(y_{j}))$. Therefore we have that:
$z_{m}=\sum_{\begin{subarray}{c}j\in\operatorname{supp}_{B}(z_{m})\\ \max(\operatorname{supp}(y_{i}))\leq k_{0}\end{subarray}}\mu_{j}y_{j}+\sum_{\begin{subarray}{c}j\in\operatorname{supp}_{B}(z_{m})\\ k_{1}\leq\min(\operatorname{supp}(y_{i}))\end{subarray}}\mu_{j}y_{j}.$
We may now express $\operatorname{supp}(z_{m})$ in two different ways:
$\bigcup_{i\in\operatorname{supp}_{A}(z_{m})}\operatorname{supp}(x_{i})$ $=\bigcup_{\begin{subarray}{c}i\in\operatorname{supp}_{A}(z_{m})\\ \max(\operatorname{supp}(x_{i}))\leq k_{0}\end{subarray}}\operatorname{supp}(x_{i})\cup\bigcup_{\begin{subarray}{c}i\in\operatorname{supp}_{A}(z_{m})\\ k_{1}\leq\min(\operatorname{supp}(x_{i}))\end{subarray}}\operatorname{supp}(x_{i}),$
$\bigcup_{j\in\operatorname{supp}_{B}(z_{m})}\operatorname{supp}(y_{j})$ $=\bigcup_{\begin{subarray}{c}j\in\operatorname{supp}_{B}(z_{m})\\ \max(\operatorname{supp}(y_{j}))\leq k_{0}\end{subarray}}\operatorname{supp}(y_{j})\cup\bigcup_{\begin{subarray}{c}j\in\operatorname{supp}_{B}(z_{m})\\ k_{1}\leq\min(\operatorname{supp}(y_{j}))\end{subarray}}\operatorname{supp}(y_{j}).$
This implies that:
$\bigcup_{\begin{subarray}{c}i\in\operatorname{supp}_{A}(z_{m})\\ \max(\operatorname{supp}(x_{i}))\leq k_{0}\end{subarray}}\operatorname{supp}(x_{i})$ $=\bigcup_{\begin{subarray}{c}j\in\operatorname{supp}_{B}(z_{m})\\ \max(\operatorname{supp}(y_{j}))\leq k_{0}\end{subarray}}\operatorname{supp}(y_{j}),$
$\bigcup_{\begin{subarray}{c}i\in\operatorname{supp}_{A}(z_{m})\\ k_{1}\leq\min(\operatorname{supp}(x_{i}))\end{subarray}}\operatorname{supp}(x_{i})$ $=\bigcup_{\begin{subarray}{c}j\in\operatorname{supp}_{B}(z_{m})\\ k_{1}\leq\min(\operatorname{supp}(y_{j}))\end{subarray}}\operatorname{supp}(y_{j}).$
Therefore, if we let:
$w^{\prime}:=\sum_{\begin{subarray}{c}j\in\operatorname{supp}_{B}(z_{m})\\ \max(\operatorname{supp}(y_{j}))\leq k_{0}\end{subarray}}\mu_{j}y_{j},$
then we have that $\operatorname{supp}(w)=\operatorname{supp}(w^{\prime})$. Furthermore, for any $k\in\operatorname{supp}(w)$, $w(k)=z_{m}(k)=w^{\prime}(k)$, so $w=w^{\prime}$. It’s clear from the definition that $w\in\langle A\rangle$ and $w^{\prime}\in\langle B\rangle$, so we have that $w\in\langle A\rangle\cap\langle B\rangle$, as desired. ∎
We now choose an increasing sequence $(n_{i})_{i<\omega}$ of representatives of all equivalence classes in $\sim$ (i.e. $\bigcup_{i<\omega}[n_{i}]_{\sim}=K$). We let $C:=(z_{n_{i}})_{i<\omega}$. By Claim 3, $C$ is an infinite block sequence. We may finish the proof of this lemma by proving the following claim:
Claim 4.
$\langle C\rangle=\langle A\rangle\cap\langle B\rangle$.
Proof.
Since $z_{n_{i}}\in\langle A\rangle\cap\langle B\rangle$ for all $i$, we have that $\langle C\rangle\subseteq\langle A\rangle\cap\langle B\rangle$. Conversely, let $v\in\langle A\rangle\cap\langle B\rangle$. Let $n_{i_{0}},\dots,n_{i_{t-1}}$ enumerate all $n_{i_{j}}$’s such that $n_{i_{j}}\in\operatorname{supp}_{A}(v)$. Fix any $k_{j}\in\operatorname{supp}(x_{n_{i_{j}}})\subseteq\operatorname{supp}(z_{n_{i_{j}}})$, and let:
$\lambda_{j}:=\frac{v(k_{j})}{z_{n_{i_{j}}}(k_{j})}.$
Note that $\lambda_{j}\neq 0$ for all $j$. Now let:
$v^{\prime}:=\sum_{j<t}\lambda_{j}z_{n_{i_{j}}}.$
Clearly $v^{\prime}\in\langle C\rangle$. We shall show that $v=v^{\prime}$, which proves the claim.
We first note that $\operatorname{supp}(v^{\prime})=\operatorname{supp}(v)$ – observe that $\operatorname{supp}_{A}(v^{\prime})=\bigcup_{j<t}\operatorname{supp}(z_{n_{i_{j}}}$). If $m\in\operatorname{supp}_{A}(v)$, then $m]_{\sim}\subseteq\operatorname{supp}_{A}(v)$ and $m]_{\sim}=[n_{i_{j}}]_{\sim}=\operatorname{supp}_{A}(z_{n_{i_{j}}})$ for some $j$, so $m\in\operatorname{supp}_{A}(v^{\prime})$. Conversely, if $m\in\operatorname{supp}(z_{n_{i_{j}}})$, then since $n_{i_{j}}\in\operatorname{supp}_{A}(v)$, we have that $m]_{\sim}=[n_{i_{j}}]_{\sim}\subseteq\operatorname{supp}_{A}(v)$, so $m\in\operatorname{supp}(v)$.
It remains to show that $v^{\prime}(k)=v(k)$ for all $k\in\operatorname{supp}(v)$. We fix some $j<t$, and observe that:
$v^{\prime}(k_{j})=\frac{v(k_{j})}{z_{n_{i_{j}}}(k_{j})}\cdot z_{n_{i_{j}}}(k_{j})=v(k_{j}),$
Since $v\in\langle A\rangle$ and $k_{j}\in\operatorname{supp}(x_{n_{i_{j}}})\subseteq\operatorname{supp}(v)$, there exist some vectors $u,w\in\langle A\rangle$ such that $u<x_{n_{i_{j}}}<w$, and:
$v=u+\frac{v(k_{j})}{x_{n_{i_{j}}}(k_{j})}x_{n_{i_{j}}}+w.$
Similarly, there exist some vectors $u^{\prime},w^{\prime}\in\langle A\rangle$ such that $u^{\prime}<x_{n_{i_{j}}}<w^{\prime}$, and:
$v^{\prime}=u^{\prime}+\frac{v(k_{j})}{x_{n_{i_{j}}}(k_{j})}x_{n_{i_{j}}}+w^{\prime}.$
Therefore, $v-v^{\prime}\in\langle A\rangle\cap\langle B\rangle$, and $n_{i_{j}}\notin\operatorname{supp}(v-v^{\prime})$. But this implies that $n_{i_{j}}]_{\sim}\cap\operatorname{supp}(v-v^{\prime})=\emptyset$, which is only possible if for all $k\in\operatorname{supp}(z_{n_{i_{j}}})$, $v^{\prime}(k)=v(k)$. Therefore, for all $j<t$ and $k\in\operatorname{supp}(z_{n_{i_{j}}})$, $v^{\prime}(k)=v(k)$, so $v^{\prime}=v$, as desired. ∎
3. Local Ramsey theory of block subspaces
Smythe explored the local Ramsey theory of block subspaces of a countable vector space developed in [S18], mainly focusing on $\leq^{*}$-upward closed subsets of $E^{[\infty]}$ (which Smythe calls families) and their various properties. We provide an overview of this theory, along with several additional results which will prove to be useful in later sections.
3.1. Semicoideals
To avoid confusion with other uses of the word “family” such as almost disjoint families, we shall use the term “semicoideals” in place of “families”.
Definition 3.1.
A subset $\mathcal{H}\subseteq E^{[\infty]}$ is a semicoideal if it is $\leq^{*}$-upward closed. That is, if $A\in\mathcal{H}$ and $A\leq^{*}B$, then $B\in\mathcal{H}$.
To cope with the failure of the pigeonhole principle in the setting of countable vector spaces with $|\mathbb{F}|\geq 3$, Smythe introduced a weaker form of the pigeonhole property of semicoideals, which he calls fullness.
Notation 3.2.
Let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal. Given $A\in\mathcal{H}$, we write:
$\mathcal{H}\mathord{\upharpoonright}A:=\{B\in\mathcal{H}:B\leq A\}.$
Definition 3.3.
Let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal.
  1. (1)
    A subset $\mathcal{D}\subseteq E^{[\infty]}$ is $\mathcal{H}$-dense below some $A\in\mathcal{H}$ if for all $B\in\mathcal{H}\mathord{\upharpoonright}A$, there exists some $C\leq B$ such that $C\in\mathcal{D}$. A set $Y\subseteq E$ is $\mathcal{H}$-dense below $A$ if the set $\{B\in E^{[\infty]}:\langle B\rangle\subseteq Y\}$ is.
  2. (2)
    We say that $\mathcal{H}$ is full if for all $Y\subseteq E$ such that $Y$ is $\mathcal{H}$-dense below some $A\in\mathcal{H}$, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}A$ such that $\langle B\rangle\subseteq Y$.
It turns out that if $|\mathbb{F}|=2$, then fullness coincides with the local pigeonhole principle property. In this case, we say that $\mathcal{H}$ is a coideal.
Definition 3.4.
Let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal. We say that $\mathcal{H}$ is a coideal if for all $Y\subseteq E$ and $A\in\mathcal{H}$, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}A$ such that $\langle B\rangle\subseteq Y$ or $\langle B\rangle\subseteq Y^{c}$.
By Hindman’s theorem, $E^{[\infty]}$ is a coideal if $|\mathbb{F}|=2$.
Lemma 3.5.
Suppose that $|\mathbb{F}|=2$, and let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal. Then $\mathcal{H}$ is full iff $\mathcal{H}$ is a coideal.
Proof.
Suppose that $\mathcal{H}$ is a coideal. Let $Y\subseteq E$, and suppose that $Y$ is $\mathcal{H}$-dense below some $A\in\mathcal{H}$. Since $\mathcal{H}$ is a coideal, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}A$ such that $\langle B\rangle\subseteq Y$ or $\langle B\rangle\subseteq Y^{c}$. Since $\mathcal{Y}$ is $\mathcal{H}$-dense below $A$, the latter option is not possible.
Now suppose that $\mathcal{H}$ is full. Let $Y\subseteq E$, and let $A\in\mathcal{H}$. If $Y$ is $\mathcal{H}$-dense below $A$, then we are done by the fullness of $\mathcal{H}$, so assume otherwise. This means that there exists some $B\in\mathcal{H}\mathord{\upharpoonright}A$ such that for all $C\leq B$, $\langle C\rangle\not\subseteq Y$. By Hindman’s theorem, for all $C\in\mathcal{H}\mathord{\upharpoonright}B$, there exists some $D\leq C$ such that $\langle D\rangle\subseteq Y^{c}$ (as $\langle D\rangle\not\subseteq Y$). But this means that $Y^{c}$ is $\mathcal{H}$-dense below $B$, so there exists some $C\in\mathcal{H}\mathord{\upharpoonright}B$ such that $\langle C\rangle\subseteq Y^{c}$, as desired. ∎
3.2. Selectivity properties
In the process of studying mad families (particularly $\mathcal{H}(\mathcal{A})$, see Definition 4.7), Smythe extended several local Ramsey theoretic properties of coideals of $\mathbb{N}$$^{\infty}$ to which Smythe extended to semicoideals. The first such property extends the (p)-property for coideals of $\mathbb{N}$$^{\infty}$ (see Lemma 7.4, [T10]).
Definition 3.6.
Let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal.
  1. (1)
    Let $(A_{n})_{n<\omega}$ be a $\leq$-decreasing sequence in $\mathcal{H}$. We say that $B\in\mathcal{H}$ weakly diagonalises $(A_{n})_{n<\omega}$ if $B\leq A_{0}$ and $B\leq^{*}A_{n}$ for all $n$.
  2. (2)
    We say that $\mathcal{H}$ satisfies the (p)-property, or that $\mathcal{H}$ is a (p)-semicoideal, if every $\leq$-decreasing sequence in $\mathcal{H}$ has a weak diagonalisation in $\mathcal{H}$.
We shall provide a similar extension of the (q)-property (see Lemma 7.4, [T10]). Smythe also made a similar extension, which he termed as spreadness (Definition 8.11, [S18]).
Definition 3.7.
Let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal. We say that $\mathcal{H}$ satisfies the (q)-property, or that $\mathcal{H}$ is a (q)-semicoideal, if for all $A\in\mathcal{H}$ and for all increasing sequence of finite intervals $(I_{m})_{m<\omega}$, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}A$ such that for all $n$, there exists some $m$ such that:
  1. (1)
    $\langle r_{n}(B)\rangle\subseteq\langle r_{\min(I_{m})-1}(A)\rangle$, and;
  2. (2)
    $B/n\leq A/\max(I_{m})$.
In a similar vein, we extend the definition of selectivity of coideals of $\mathbb{N}$$^{\infty}$ (see Definition 7.3, [T10]) to semicoideals of block subspaces.
Definition 3.8.
For any $A\in E^{[\infty]}$ and $x\in E$, we write:
$d_{A}(x):=\min\{N<\omega:\max(\operatorname{supp}(x))<A/N\}.$
Definition 3.9.
Let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal.
  1. (1)
    Let $A\in\mathcal{H}$, and let $(A_{n})_{n<\omega}$ be a $\leq$-decreasing sequence in $\mathcal{H}$ below $A$ (i.e. $A_{0}\leq A$). We say that $B\in\mathcal{H}\mathord{\upharpoonright}A$ diagonalises $(A_{n})_{n<\omega}$ within $A$ if $B/n\leq A_{d_{A}(r_{n}(B))}$ for all $n<\omega$.
  2. (2)
    We say that $\mathcal{H}$ is selective if for all $A\in\mathcal{H}$ and $\leq$-decreasing chain $(A_{n})_{n<\omega}$ of elements in $\mathcal{H}$ below $A$, there exists a diagonalisation of $(A_{n})_{n<\omega}$ within $A$ in $\mathcal{H}$.
Lemma 3.10.
Let $\mathcal{H}\subseteq E^{[\infty]}$ be a (q)-semicoideal. Let $(A_{n})_{n<\omega}$ be a $\leq$-decreasing sequence in $\mathcal{H}$. If $B\in\mathcal{H}\mathord{\upharpoonright}A_{0}$ weakly diagonalises $(A_{n})_{n<\omega}$, then there exists some $C\in\mathcal{H}\mathord{\upharpoonright}B$ which diagonalises $(A_{n})_{n<\omega}$ within $A_{0}$.
Proof.
Let $A:=A_{0}$, and let $B\in\mathcal{H}\mathord{\upharpoonright}A$ weakly diagonalise $(A_{n})_{n<\omega}$. Let $(I_{m})_{m<\omega}$ be any increasing sequence of intervals such that $B/\max(I_{m})\leq A_{d_{A}(r_{\min(I_{m})-1}(B))}$. By the (q)-property, there exists some $C\in\mathcal{H}\mathord{\upharpoonright}B$ such that for all $n$, $\langle r_{n}(C)\rangle\subseteq\langle r_{\min(I_{m})-1}(B)\rangle$ and $C/n\leq B/\max(I_{m})$. The first property implies that $d_{A}(r_{n}(C))\leq d_{A}(r_{\min(I_{m})}(B))$, so by the second property:
$C/n\leq B/\max(I_{m})\leq A_{d_{A}(r_{\min(I_{m})}(B))}\leq A_{d_{A}(r_{n}(C))},$
so $C$ strongly diagonalises $(A_{n})_{n<\omega}$. ∎
Proposition 3.11.
Let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal. The following are equivalent:
  1. (1)
    $\mathcal{H}$ is selective.
  2. (2)
    $\mathcal{H}$ satisfies both the (p)-property and the (q)-property.
Proof.
[…]
$\langle r_{n}(B)\rangle\subseteq\langle r_{k}(A)\rangle\subseteq\langle r_{\min(I_{m(k)})}(A)\rangle,$
and:
$B/n\leq A_{d_{A}(r_{n}(B))}=A/\max(I_{m(k)}),$
as desired. ∎
We shall examine the selectivity of semicoideals generated by almost disjoint families in the next section.
4. Almost disjoint families
We recall the following:
Definition 4.1.
Two infinite-dimensional subspaces $V,W\subseteq E$ are almost disjoint if $V\cap W$ is a finite-dimensional subspace of $E$. Otherwise, we say that $V$ and $W$ are compatible.
Definition 4.2.
Let $\mathcal{A}$ be a family of infinite-dimensional subspaces of $E$. We say that $\mathcal{A}$ is almost disjoint if all subspaces in $\mathcal{A}$ are pairwise almost disjoint. We say that $\mathcal{A}$ is maximal almost disjoint (or just mad) if $\mathcal{A}$ is not strictly contained in another almost disjoint family of infinite-dimensional subspaces.
Since every infinite-dimensional subspace of $E$ contains a block subspace, an almost disjoint family of subspaces $\mathcal{A}$ is mad iff there is no block subspace which is almost disjoint from every element of $\mathcal{A}$.
We dedicate this section to some basic definitions/notations and results of almost disjoint families of block subspaces, and their relationship with the local Ramsey theory of semicoideals of block subspaces.
4.1. Uncountability of almost disjoint families
We begin with the most fundamental question of the theory of mad families – the uncountability of infinite mad families. Consider the following two cardinals:
[…]
It is clear that $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}\leq\mathfrak{c}$. By Proposition 2.5 of [S19], $\aleph_{1}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ (in fact, $\operatorname{non}(\mathcal{M})\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$), so $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ is a cardinal invariant. One may ask if the same holds for $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ - that is, if $\aleph_{1}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ is true. Smythe’s proof of $\aleph_{1}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ heavily relies on the following lemma:
Lemma 4.3 (Lemma 2.3, [S18]).
Let $A\in E^{[\infty]}$ be a block subspace, and let $a\in E^{[<\infty]}$ be a finite-dimensional block subspace. Then there exists some $N<\omega$, that depends only on $A$ and $\max(\operatorname{supp}(a))$, such that for all $x>N$ such that $x\notin\langle A\rangle$,
$\langle a^{\frown}x\rangle\cap\langle A\rangle=\langle a\rangle\cap\langle A\rangle.$
He remarked that removing the “block” requirement from this lemma seems difficult. $\aleph_{1}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ was also stated as Proposition 3.1 of [K00], but Kolman further assumed that in the countable almost disjoint family of subspaces $\{V_{n}:n<\omega\}$, $V_{n}\cap\sum_{i\in I}V_{i}$ is finite-dimensional for all finite $n\notin I\subseteq\omega$. This is generally not satisfied by families of subspaces that are pairwise almost disjoint. The author raised this question on MathOverflow, of which user527492 provided an affirmative answer (see [u25]).
Lemma 4.4.
$\aleph_{1}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$.
Proof.
Let $\{V_{n}:n<\omega\}$ be a countable family of almost disjoint subspaces of $E$, and we wish to find some infinite-dimensional subspace $W$ which is almost disjoint from $V_{n}$ for all $n$. We shall first prove a useful claim.
Claim 5.
Let $U,V,W\subseteq E$ be subspaces such that $\dim(V\cap W)<\infty$ and $\dim(U)<\infty$. Then $\dim((U+V)\cap W)<\infty$.
Proof.
Define a linear map $\phi:V\times W\to E$ by stipulating that $\phi(v,w):=w-v$. Let $\rho:E\to E/U$ be the canonical quotient map, and let $\psi:=\rho\circ\pi$. Note that:
$\dim(\ker(\psi))\leq\dim(\ker(\phi))+\dim(\ker(\rho))<\infty.$
Let $\pi:V\oplus W\to W$ be the projection map. Observe that for all $w\in W$, we have that:
$\phantom{{}\iff{}}w\in(U+V)\cap W$
$\iff w=u+v\text{ for some $u\in U$, $v\in V$}$
$\iff w-v\in U\text{ for some $v\in V$}$
$\iff(v,w)\in\ker(\psi)\text{ for some $v\in V$}.$
This implies that $\pi$ maps $\ker(\psi)$ surjectively onto $(U+V)\cap W$, and since $\ker(\psi)$ is of finite dimension, so is $(U+V)\cap W$. ∎
We shall define the subspace $W=\{x_{n}:n<\omega\}$ inductively as follows: Let $x_{0}\in V_{0}$ be any vector. Now assume that we have defined a linearly independent set $\{x_{i}:i<n\}$ such that for all $k<n$, we have that:
$\operatorname{span}\{x_{i}:i<n\}\cap V_{k}\subseteq\operatorname{span}\{x_{i}:i\leq k\}.$
We wish to find some vector $x_{n}\in V_{n}$ such that for all $k<n$, we have that:
$\operatorname{span}\{x_{i}:i\leq n\}\cap V_{k}\subseteq\operatorname{span}\{x_{i}:i\leq k\}.$
If we can complete this induction step, then we are done as we have that $W\cap V_{n}\subseteq\operatorname{span}\{x_{0},\dots,x_{n}\}$ is finite-dimensional for all $n$.
For each $k<n$, let $U_{k}:=V_{k}+\operatorname{span}\{x_{i}:i<n\}$. By Claim 5, $V_{n}\cap U_{k}$ is of finite dimension. Since every infinite-dimensional vector space is not a finite union of finite-dimensional subspaces, we have that:
$V_{n}\setminus\bigcup_{k<n}V_{n}\cap U_{k}\neq\emptyset,$
so we may let $x_{n}$ be any vector in this set. We shall show that this vector $x_{n}$ works. We first note that $\{x_{i}:i\leq n\}$ is linearly independent, as $\operatorname{span}\{x_{i}:i<n\}\subseteq U_{0}$ and $x_{n}\notin U_{0}$. Now let $k<n$ and $x\in\operatorname{span}\{x_{i}:i<n+1\}\cap V_{k}$, so we have that $x=\sum_{i\leq n}\lambda_{i}x_{i}\in V_{k}$. If $\lambda_{n}\neq 0$, then we have that:
$x_{n}=\lambda_{n}^{-1}\left(x-\sum_{i<n}\lambda_{i}x_{i}\right)\in V_{n}\cap U_{k},$
contradicting the choice of $x_{n}$. Therefore, $\lambda_{n}=0$, which implies that:
$x\in\operatorname{span}\{x_{i}:i<n\}\cap V_{k}=\operatorname{span}\{x_{i}:i\leq k\}\cap V_{k},$
as desired. ∎
Smythe asked in [S19] whether $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}=\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ holds in $\mathsf{ZFC}$, and this problem remains open to the author’s knowledge. In the same paper, Smythe showed that every finite almost disjoint family of block subspaces of cardinality greater than $1$ is not maximal. We extend this fact to arbitrary subspaces.
Proposition 4.5.
Let $\mathcal{A}$ be a finite almost disjoint family of infinite-dimensional subspaces of $E$ such that $|\mathcal{A}|>1$. Then $\mathcal{A}$ is not maximal.
Proof.
Let $\mathcal{A}=\{V_{k}:k<n\}$ be a finite almost disjoint family of infinite-dimensional subspaces of $E$.
Claim 6.
$E\neq\bigcup_{k<n}V_{k}$.
Proof.
We first assume that for all $j\neq k$, $V_{j}\cap V_{k}=\{0\}$. If $\mathbb{F}$ is infinite, then we are done, as any subspace over an infinte field is not a finite union of proper subspaces. Assume that $|\mathbb{F}|=N<\infty$, and suppose for a contradiction that $E=\bigcup_{k<n}V_{k}$. Let $M$ be some large number (which we shall fix later), and let $W:=\operatorname{span}\{e_{n}:n<M\}$. Let $U_{k}:=W\cap V_{k}$. Note that $W=\bigcup_{j<n}U_{k}$.
Observe that $\dim(W)=M$, so if $U\subseteq W$ is a subspace such that $\dim(U)>\frac{M}{2}$, then $U$ intersects every proper subspace of $M$ non-trivially. Since $\{V_{k}:k<n\}$ is a collection of proper subspaces of $M$ with pairwise trivial intersections with at least $2$ elements, each $V_{k}$ is at most $\frac{M}{2}$-dimensional. Consequently, every $U_{i}$ contains at most $N^{\frac{M}{2}}$ vectors, which implies that:
$\left|\bigcup_{k<n}V_{k}\right|\leq n\cdot N^{\frac{M}{2}}.$
On the other hand, $|W|=N^{M}$. Therefore, we may choose $M$ large enough so that $N^{M}>n\cdot N^{\frac{M}{2}}$, contradicting the fact that $W=\bigcup_{k<n}U_{k}$.
For the general case, suppose that $\{V_{k}:k<n\}$ is almost disjoint and $E=\bigcup_{k<n}V_{k}$. Let $M$ be large enough so that $M>\max(\operatorname{supp}[V_{j}\cap V_{k}])$ for all $j\neq k$. Let $W^{\prime}:=\operatorname{span}\{e_{n}:n\geq M\}$, and we have that $W^{\prime}=\bigcup_{k<n}W^{\prime}\cap V_{k}$. Then $\{W^{\prime}\cap V_{k}:k<n\}$ is a collection of infinite-dimensional subspaces of $W^{\prime}$ with pairwise trivial intersections, contradicting the argument above. ∎
We show now construct some $V=\operatorname{span}\{x_{m}:m<\omega\}\subseteq E$ which intersects $V_{k}$ trivially for every $k<n$. Let $x_{0}\in E\setminus\bigcup_{k<n}V_{k}$ be any vector. Now suppose that we have defined $x_{i}$ for $i<m$ such that for all $k<n$:
$\operatorname{span}\{x_{i}:i<m\}\cap V_{k}=\{0\}.$
For each $k<n$, let $U_{k}:=V_{k}+\operatorname{span}\{x_{i}:i<m\}$. By the claim in the proof of Lemma 4.4, $\{U_{k}:k<n\}$ is an almost disjoint family of subspaces of $E$, so by applying Claim 6 to $\{U_{k}:k<n\}$, $E\neq\bigcup_{k<n}U_{k}$. Let $x_{m}\in E\setminus\bigcup_{k<n}U_{k}$. Note that $\{x_{i}:i\leq m\}$ is linearly independent, as $x_{i}\in U_{0}$ for all $i<m$ but $x_{m}\notin U_{0}$. We shall show that:
$\operatorname{span}\{x_{i}:i\leq m\}\cap V_{k}=\{0\}.$
If we can complete this induction step, then we are done as $V\cap V_{k}=\bigcup_{m<\omega}\operatorname{span}\{x_{i}:i<m\}\cap V_{k}=\{0\}$ for all $k<n$.
Let $y\in\operatorname{span}\{x_{i}:i\leq m\}\cap V_{k}$ for some $k<n$. Then $y=\sum_{i\leq m}\lambda_{i}x_{i}$ for some $\lambda_{i}\in\mathbb{F}$. If $\lambda_{m}\neq 0$, then we have that:
$x_{m}=\lambda_{m}^{-1}\left(y-\sum_{i<m}\lambda_{i}x_{i}\right)\in U_{k},$
contradicting the choice of $x_{m}$. Therefore, $\lambda_{m}=0$, which implies that:
$y\in\operatorname{span}\{x_{i}:i<m\}\cap V_{k}=\{0\},$
as desired. ∎
4.2. Ideals and semicoideals generated by almost disjoint families
If $\mathcal{A}$ is an almost disjoint family of $\omega]^{\omega}$, then the following families of sets are often studied:
$\mathcal{I}(\mathcal{A})$ $:=\{Y\subseteq\omega:Y\subseteq^{*}A_{0}\cup\cdots\cup A_{n-1}\text{ for some $A_{0},\dots,A_{n-1}\in\mathcal{A}$}\},$
$\mathcal{I}^{+}(\mathcal{A})$ $:=\mathcal{P}(\omega)\setminus\mathcal{I}(\mathcal{A}),$
$\mathcal{I}^{++}(\mathcal{A})$ $:=\{Y\subseteq\omega:\{A\in\mathcal{A}:Y\cap A\text{ is infinite}\}\text{ is infinite}\},$
$\mathcal{H}(\mathcal{A})$ $:=\mathcal{I}^{++}(\mathcal{A}).$
It is simple to verify that $\mathcal{I}^{+}(\mathcal{A})=\mathcal{I}^{++}(\mathcal{A})$ iff $\mathcal{A}$ is maximal. Motivated by this, Smythe introduced a version of $\mathcal{H}(\mathcal{A})$ for almost disjoint families of block subspaces. $\mathcal{H}(\mathcal{A})$ is a semicoideal, and Smythe studied the conditions that $\mathcal{A}$ must satisfy to make $\mathcal{H}(\mathcal{A})$ a full semicoideal.
Definition 4.6.
Let $Y\subseteq E$.
  1. (1)
    $Y$ is big if there exists some $A\in E^{[\infty]}$ such that $\langle A\rangle\subseteq Y$.
  2. (2)
    $Y$ is small if it is not big.
  3. (3)
    $Y$ is very small if for all $Z\subseteq E$ such that $Z$ is small, $Y\cup Z$ is small.
Definition 4.7.
Let $\mathcal{A}$ be an almost disjoint family of infinite-dimensional subspaces. We then define the following:
$\mathcal{I}(\mathcal{A})$ $:=\left\{Y\subseteq E:\begin{matrix}Y\setminus(\langle A_{0}\rangle\cup\cdots\cup\langle A_{n-1}\rangle)\text{ is small}\\ \text{for some }A_{0},\dots,A_{n-1}\in\mathcal{A}\end{matrix}\right\},$
$\mathcal{I}^{+}(\mathcal{A})$ $:=\mathcal{P}(E)\setminus\mathcal{I}(\mathcal{A}),$
$\mathcal{I}^{++}(\mathcal{A})$ $:=\{Y\subseteq E:\{A\in\mathcal{A}:Y\cap\langle A\rangle\text{ is big}\}\text{ is infinite}\},$
$\mathcal{H}^{-}(\mathcal{A})$ $:=\{B\in E^{[\infty]}:\langle B\rangle\in\mathcal{I}^{+}(\mathcal{A})\},$
$\mathcal{H}(\mathcal{A})$ $:=\{B\in E^{[\infty]}:\langle B\rangle\in\mathcal{I}^{++}(\mathcal{A})\}.$
Lemma 4.8.
Let $\mathcal{A}\subseteq E^{[\infty]}$ be an infinite almost disjoint family, and let $Y\subseteq E$. The following are equivalent:
  1. (1)
    $\langle Y\rangle\in\mathcal{I}^{+}(\mathcal{A})$.
  2. (2)
    $\langle Y\rangle\in\mathcal{I}^{++}(\mathcal{A})$ or there exists some $B\in E^{[\infty]}$, with $\langle B\rangle\subseteq\langle Y\rangle$, such that $B$ is almost disjoint from every element of $\mathcal{A}$.
Consequently, if $\mathcal{A}$ is mad, then $\mathcal{H}^{-}(\mathcal{A})=\mathcal{H}(\mathcal{A})$.
Proof.
We remark that for any infinite almost disjoint family $\mathcal{A}$, $\mathcal{I}^{++}(\mathcal{A})\subseteq\mathcal{I}^{+}(\mathcal{A})$.
[…]
We conclude this part with three useful lemmas we shall use repeatedly in this paper.
Lemma 4.9.
Let $A\in E^{[\infty]}$ and $Y\subseteq E$. Then $\langle A\rangle\cap\langle Y\rangle$ is small iff $\langle A/N\rangle\cap\langle Y\rangle=\emptyset$ for some $N$. Consequently:
  1. (1)
    If $A,B\in E^{[\infty]}$, then $A$ and $B$ are almost disjoint iff $\langle A/N\rangle\cap\langle B\rangle=\emptyset$ for some $N$.
  2. (2)
    If $A/N$ and $B$ are almost disjoint for some $N$, then $A$ and $B$ are almost disjoint.
Proof.
Suppose that $\langle A\rangle\cap\langle Y\rangle$ is big. Let $B\in E^{[\infty]}$ such that $\langle B\rangle\subseteq\langle A\rangle\cap\langle Y\rangle$. For any $N$, there exists some $M$ such that $B/M\leq A/N$. This implies that $\langle B/M\rangle\subseteq\langle A/N\rangle\cap\langle Y\rangle$, so in particular, $\langle A/N\rangle\cap\langle Y\rangle\neq\emptyset$ for all $N$.
Conversely, suppose that $\langle A/N\rangle\cap\langle Y\rangle\neq\emptyset$ for all $N$. We construct some $B=(y_{n})_{n<\omega}\in E^{[\infty]}$ such that $\langle B\rangle\subseteq\langle A\rangle\cap\langle Y\rangle$ inductively as follows: Assuming that $y_{0},\dots,y_{n-1}$ have been defined so far, let $N$ be large enough so that $\operatorname{supp}(y_{n-1})<A/N$. We let $y_{n}\in\langle A/N\rangle\cap\langle Y\rangle$ be any element. Then $\langle B\rangle\subseteq\langle A\rangle\cap\langle Y\rangle$, so $\langle A\rangle\cap\langle Y\rangle$ is big. ∎
Lemma 4.10.
Let $A\in E^{[\infty]}$, and suppose that $\{B_{n}:n<\omega\}\subseteq E^{[\infty]}$ is such that $B_{n}\leq A$ for all $n$. Then there exists some $B\leq A$ such that $B$ is compatible with $B_{n}$ for all $n$, and:
$\langle B\rangle\subseteq\left\langle\bigcup_{n<\omega}\langle B_{n}\rangle\right\rangle\subseteq\langle A\rangle.$ (*)
Proof.
Let $\Gamma:\omega\to\omega$ be any infinite-to-one surjection, i.e. a surjection such that $\Gamma^{-1}(m)$ is infinite for all $m<\omega$. We define $B=(x_{n})_{n<\omega}$ inductively as follows: If $(x_{k})_{k<n}$ has been defined, let $x_{n}\in\langle B_{\Gamma(n)}\rangle$ be any element such that $(x_{k})_{k<n}<x_{n}$. Then $(x_{n})_{n\in\Gamma^{-1}(m)}$ witnesses that $B$ and $B_{m}$ are compatible for all $m<\omega$, and since $x_{n}\in\bigcup_{m<\omega}\langle B_{m}\rangle$ for all $n$, the identity (*) is satisfied. ∎
Lemma 4.11.
Let $\mathcal{A}\subseteq E^{[\infty]}$ be an almost disjoint family, and let $Y\subseteq E$. Suppose further that $\langle Y\rangle\in\mathcal{I}^{+}(\mathcal{A})$ and there does not exist any $B\in E^{[\infty]}$ such that $\langle B\rangle\subseteq\langle Y\rangle$ and $B$ is almost disjoint from every element of $\mathcal{A}$. Then there exists an infinite set $\{A_{n}:n<\omega\}\subseteq\mathcal{A}$ and an infinite disjoint family $\{B_{n}:n<\omega\}$ such that for all $n$:
  1. (1)
    $B_{n}\leq A_{n}$ for some $A_{n}\in\mathcal{A}$.
  2. (2)
    $\langle B_{n}\rangle\subseteq\langle Y\rangle$.
Note that these two properties imply that $\langle Y\rangle\in\mathcal{I}^{++}(\mathcal{A})$.
Proof.
Since $\langle Y\rangle\in\mathcal{I}^{+}(\mathcal{A})$, $\langle Y\rangle$ is big, so there exists some $B_{0}^{\prime}$ such that $\langle B_{0}^{\prime}\rangle\subseteq\langle Y\rangle$. By our assumption, $B_{0}^{\prime}$ is compatible with some $A_{0}\in\mathcal{A}$. Let $B_{0}$ be such that $\langle B_{0}\rangle\subseteq\langle B_{0}^{\prime}\rangle\cap\langle A_{0}\rangle$. Since $\langle Y\rangle\in\mathcal{I}^{+}(\mathcal{A})$, $\langle Y\rangle\setminus\langle A_{0}\rangle$ is big. Given an ordinal $\alpha$, suppose that we have defined an infinite disjoint family $\{B_{\beta}:\beta<\alpha\}$ and $\{A_{\beta}:\beta<\alpha\}\subseteq\mathcal{A}$ such that $B_{\beta}\leq A_{\beta}$ for all $\beta<\alpha$, and $\langle B_{\beta}\rangle\subseteq\langle Y\rangle$ for all $\beta<\alpha$. If $Y\setminus\bigcup_{\beta<\alpha}\langle B_{\beta}\rangle$ is small, we terminate the construction. Otherwise, we let $B_{\alpha}^{\prime}$ be such that $\langle B_{\alpha}^{\prime}\rangle\subseteq Y\setminus\bigcup_{\beta<\alpha}\langle B_{\beta}\rangle$, and let $A_{\alpha}\in\mathcal{A}$ be compatible with $B_{\alpha}^{\prime}$. We then let $B_{\alpha}$ be such that $\langle B_{\alpha}\rangle\subseteq\langle B_{\alpha}^{\prime}\rangle\cap\langle A_{\alpha}\rangle$. Since $\langle Y\rangle$ is countable and the $B_{\alpha}$’s are disjoint, this induction must terminate at some countable stage $\alpha$, and since $\langle Y\rangle\in\mathcal{I}^{+}(\mathcal{A})$, $\{A_{\beta}:\beta<\alpha\}$ must be infinite. ∎
4.3. Selectivity properties revisited
We now return to studying the combinatorial properties of the semicoideals $\mathcal{H}^{-}(\mathcal{A})$ and $\mathcal{H}(\mathcal{A})$. In Mathias’ process of proving that there are no analytic mad families of $\mathbb{N}$$^{\infty}$, he proved that if $\mathcal{A}$ is an almost disjoint family in $\mathbb{N}$$^{\infty}$, then $\mathcal{H}^{-}(\mathcal{A})$ is a selective coideal. In [S19], Smythe proved that if $\mathcal{A}$ is a mad family in $(E^{[\infty]},\leq,r)$, then $\mathcal{H}(\mathcal{A})$ is a (p)-semicoideal. We prove a few extensions of this result.
Proposition 4.12.
Let $\mathcal{A}\subseteq E^{[\infty]}$ be an almost disjoint family. Then $\mathcal{H}^{-}(\mathcal{A})$ is a (p)-semicoideal.
Proof.
Let $(B_{n})_{n<\omega}$ be a $\leq$-decreasing sequence in $\mathcal{H}^{-}(\mathcal{A})$. If there exists some $C$ which weakly diagonalises $(B_{n})_{n<\omega}$ and $C$ is almost disjoint from every element of $\mathcal{A}$, then $C\in\mathcal{H}^{-}(\mathcal{A})$ and we are done, so assume otherwise. Let $C_{0}$ be a weak diagonalisation of $(B_{n})_{n<\omega}$. By our assumption, $C_{0}$ is compatible with some $A_{0}\in\mathcal{A}$, so by taking a further block subspace of $C_{0}$ if necessary, we may assume that $C_{0}\leq A_{0}$. Now assume, for the induction hypothesis, that we have defined $C_{k},A_{k}$ for $k<m$ such that:
  1. (1)
    $C_{k}\leq B_{k}$ is a diagonalisation of $(B_{n})_{n<\omega}$.
  2. (2)
    $C_{k}\leq A_{k}$ and $A_{k}\in\mathcal{A}$.
  3. (3)
    If $i,j<m$ then $A_{i}\neq A_{j}$.
Claim 7.
There exists a $\leq$-decreasing sequence $(B_{n}^{m})_{n<\omega}$ such that for all $n$:
  1. (1)
    $B_{n}^{m}\leq B_{n}$, and;
  2. (2)
    $B_{n}^{m}$ is almost disjoint from $A_{k}$ for all $k<m$.
Proof.
For any $n$, since $B_{n}\in\mathcal{H}^{-}(\mathcal{A})$ there exists some $D_{n}\leq B_{n}$ that is disjoint from $A_{k}$ for all $k<m$. Let $\Gamma:\omega\to\omega$ be any infinite-to-one surjection. We define $D=(x_{n})_{n<\omega}$ as follows: Assuming that $(x_{i})_{i<n}$ has been defined so far, by Lemma 4.3 we may let $N$ be large enough so that for all $k<m$ and $x\in\langle B_{0}/N\rangle$ such that $x\notin\langle A_{k}\rangle$, we have that:
$\langle x_{0},\dots,x_{n-1},x\rangle\cap\langle A_{k}\rangle=\langle x_{0},\dots,x_{n-1}\rangle\cap\langle A_{k}\rangle$
We then let $x_{n}\in\langle D_{\Gamma(n)}\rangle\cap\langle B_{0}/N\rangle$ be any element. $D$ is almost disjoint from $A_{k}$ for all $k<m$, as:
$\langle D\rangle\cap\langle A_{k}\rangle$ $=\bigcup_{i<\omega}\langle x_{0},\dots,x_{i}\rangle\cap\langle A_{k}\rangle=\langle x_{0},\dots,x_{k-1}\rangle\cap\langle A_{k}\rangle.$
Let $B_{n}^{m}:=(x_{i})_{i<\omega,\Gamma(i)\geq n}$ Then $B_{n}^{m}\leq B_{n}$, $(B_{n}^{m})_{n<\omega}$ is a $\leq$-decreasing sequence, and $B_{n}^{m}\leq D$, so $B_{n}^{m}$ is almost disjoint from $A_{k}$ for all $k\leq m$. ∎
Let $C_{m}\leq B_{m}^{m}$ be a weak diagonalisation of $(B_{n}^{m})_{n<\omega}$. By our assumption at the beginning, there exists some $A_{m}\in\mathcal{A}$ which is compatible with $C_{m}$, and we may assume that $C_{m}\leq A_{m}$. Since $C_{m}\leq B_{m}^{m}$, and $B_{m}^{m}$ is almost disjoint from $A_{k}$ for all $k<m$, we have that $A_{k}\neq A_{m}$ for all $k<m$. Having completed the induction, we let $C$ be a weak diagonalisation of $(C_{m})_{m<\omega}$. Then $C\leq^{*}C_{0}\leq^{*}B_{n}$ for any $n$, so $C$ weak diagonalises $(B_{n})_{n<\omega}$. Since $C\leq^{*}C_{m}\leq A_{m}$ for any $m$, $C$ is compatible with infinitely many elements of $\mathcal{A}$, so $C\in\mathcal{H}^{-}(\mathcal{A})$. ∎
Proposition 4.13.
Let $\mathcal{A}\subseteq E^{[\infty]}$ be an almost disjoint family. Then $\mathcal{H}^{-}(\mathcal{A})$ is a selective semicoideal.
Proof.
By Proposition 4.12, it remains to show that $\mathcal{H}^{-}(\mathcal{A})$ is a (q)-semicoideal. Let $A\in\mathcal{H}^{-}(\mathcal{A})$, and let $(I_{m})_{m<\omega}$ be an increasing sequence of finite intervals. We split our proof into two cases.
If there is some $B\leq A$ which is almost disjoint from every element of $\mathcal{A}$, then $C\in\mathcal{H}^{-}(\mathcal{A})$ for all $C\leq B$, so any “sufficiently sparse” subsequence of $B$ would do the trick. More precisely, if $B=(y_{n})_{n<\omega}$, then we define a subsequence of integers $(n_{k})_{k<\omega}$ as follows: Let $n_{0}:=0$, and if $n_{k}$ is defined, let $m$ be the least integer such that $n_{k}<\min(I_{m})$. We then define $n_{k+1}:=\max(I_{m})+1$, and $(y_{n_{k}})_{k<\omega}\in\mathcal{H}^{-}(\mathcal{A})$ satisfies the conclusion of the (q)-property for $A$ and $(I_{m})_{m<\omega}$.
Suppose there is no $B\leq A$ which is almost disjoint from every element of $\mathcal{A}$. By Lemma 4.11, there exists an infinite $\{A_{n}:n<\omega\}\subseteq\mathcal{A}$ and an infinite disjoint family $\{B_{n}:n<\omega\}$ such that $\langle B_{n}\rangle\subseteq\langle A\rangle\cap\langle A_{n}\rangle$ for all $n$. Let $\Gamma:\omega\to\omega$ be any infinite-to-one surjection. We write $A=(x_{n})_{n<\omega}$, and define $B=(y_{n})_{n<\omega}$ as follows: If $(y_{k})_{k<n}$ has been defined, we let $m$ be the least integer such that $\langle r_{n}(B)\rangle\subseteq\langle x_{0},\dots,x_{\min(I_{m})-1}\rangle$. We let $y_{n}\in\langle B_{\Gamma(n)}/\max(I_{m})\rangle$ be any element. Then $(y_{n})_{n\in\Gamma^{-1}(m)}$ witnesses that $B$ and $B_{m}$ are compatible for all $m<\omega$, and since $y_{n}\in\langle A\rangle$ for all $n$, $B\leq A$ and $B\in\mathcal{H}^{-}(\mathcal{A})$. ∎
5. Construction of mad families of subspaces
Definition 5.1.
A mad family $\mathcal{A}\subseteq E^{[\infty]}$ is full if $\mathcal{H}(\mathcal{A})$ is full.
We dedicate this section to the proof of Theorem 1.5.
5.1. Block-splitting families
We introduce/recall the cardinal invariants which are relevant to our proof of Theorem 1.5. If $f,g\in\omega^{\omega}$, we write $f\leq^{*}g$ if $f(n)\leq g(n)$ for almost all $n<\omega$. We say that $\mathcal{F}\subseteq\omega^{\omega}$ is bounded if there exists some $g\in\omega^{\omega}$ such that $f\leq^{*}g$ for all $f\in\mathcal{F}$, and unbounded otherwise. The bounding number $\mathfrak{b}$ is defined as:
$\mathfrak{b}:=\{|\mathcal{F}|:\mathcal{F}\subseteq\omega^{\omega}\text{ is unbounded}\}.$
If $S\in[\omega]^{\omega}$ and $A\in[\omega]^{\omega}$, we say that $S$ splits $A$ if $A\cap S$ and $A\setminus S$ are both infinite. A splitting family is a family $\mathcal{S}\subseteq[\omega]^{\omega}$ such that for all $A\in[\omega]^{\omega}$, there exists some $S\in\mathcal{S}$ which splits $A$. The splitting number is defined as:
$\mathfrak{s}:=\{|\mathcal{S}|:\mathcal{S}\subseteq[\omega]^{\omega}\text{ is a splitting family}\}.$
It turns out that both of these cardinals are closely related to block-splitting families. If $\mathcal{P}=\{P_{n}:n<\omega\}$ is an interval partition, we say that a set $S\in[\omega]^{\omega}$ splits $\mathcal{P}$ if $\{n<\omega:P_{n}\subseteq S\}$ and $\{n<\omega:P_{n}\cap S=\emptyset\}$ are both infinite. A block-splitting family is a family $\mathcal{S}\subseteq[\omega]^{\omega}$ such that every interval partition is block-split by some $S\in\mathcal{S}$. The following lemma is due to Kamburelis-Weglorz.
Lemma 5.2 ([KW96]).
$\max\{\mathfrak{b},\mathfrak{s}\}=\min\{|\mathcal{S}|:\mathcal{S}\text{ is a block-splitting family}\}$.
We abbreviate $\mathfrak{bs}=\max\{\mathfrak{b},\mathfrak{s}\}$. By Corollary 2.11 of [S19], the identity $\operatorname{non}(\mathcal{M})\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ holds for any countable field $\mathbb{F}$. Since $\mathfrak{b}\leq\operatorname{non}(\mathcal{M})$ and $\mathfrak{s}\leq\operatorname{non}(\mathcal{M})$ (see, for instance, [handbook]), we may conclude that:
Lemma 5.3.
$\mathfrak{bs}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$.
5.2. Proof of Theorem 1.5
Lemma 5.4.
Let $A,B\in E^{[\infty]}$.
  1. (1)
    If $\langle A\rangle\cap\langle B\rangle$ is small, then $\langle A\rangle\cap\langle B\rangle$ is very small.
  2. (2)
    If $\langle A\rangle\setminus\langle B\rangle$ is small, then $\langle A\rangle\setminus\langle B\rangle$ is very small.
Consequently, it is not possible to have both $\langle A\rangle\cap\langle B\rangle$ and $\langle A\rangle\setminus\langle B\rangle$ small.
Proof.
Assume that $\langle A\rangle\cap\langle B\rangle$ is small. Let $Y\subseteq E$, and suppose that $\langle C\rangle\subseteq(\langle A\rangle\cap\langle B\rangle)\cup Y$ for some $C=(x_{n})_{n<\omega}\in E^{[\infty]}$. Since $\langle A\rangle\cap\langle B\rangle$ is a finite-dimensional space, we have that $M:=\max(\operatorname{supp}[\langle A\rangle\cap\langle B\rangle])<\infty$. Let $N$ be large enough so that $\operatorname{supp}(x_{n})>M$. Then $\langle C/N\rangle\cap(\langle A\rangle\cap\langle B\rangle)=\emptyset$, so $\langle C/N\rangle\subseteq Y$. Therefore, $Y$ is big.
Now assume that $\langle A\rangle\setminus\langle B\rangle$ is small. Let $Y\subseteq E$, and suppose that $\langle C\rangle\subseteq(\langle A\rangle\setminus\langle B\rangle)\cup Y$. We consider three cases. For the first case, suppose that $C$ is almost disjoint from $A$. By Lemma 4.9, let $N$ be large enough so that $\langle C/N\rangle\cap\langle A\rangle=\emptyset$. Then this implies that $\langle C/N\rangle\subseteq Y$, so $Y$ is big.
For the second case, suppose that $C$ is compatible with $B$. Let $D\in E^{[\infty]}$ be such that $\langle D\rangle\subseteq\langle C\rangle\cap\langle B\rangle$. Then $\langle D\rangle\subseteq(\langle A\rangle\setminus\langle B\rangle)\cup Y$, and since $\langle D\rangle\cap(\langle A\rangle\setminus\langle B\rangle)=\emptyset$, we have that $\langle D\rangle\subseteq Y$. Therefore, $Y$ is big.
For the last case, suppose that $C$ is compatible with $A$ but almost disjoint from $B$. Let $D\in E^{[\infty]}$ be such that $\langle D\rangle\subseteq\langle C\rangle\cap\langle A\rangle$, and note that $D$ is also almost disjoint from $B$. By Lemma 4.9, let $N$ be large enough so that $\langle D/N\rangle\cap\langle B\rangle=\emptyset$. Since $D/N\leq A$, this implies that $\langle D/N\rangle\subseteq\langle A\rangle\setminus\langle B\rangle$, contradicting our assumption that $\langle A\rangle\setminus\langle B\rangle$ is small. ∎
Lemma 5.5.
There exists a collection $\{Z_{\alpha}:\alpha<\mathfrak{bs}\}$ of subsets of $E$ such that for all almost disjoint families $\mathcal{A}\subseteq E^{[\infty]}$ and $B\in\mathcal{H}^{-}(\mathcal{A})$, the following properties hold:
  1. (i)
    For all $\alpha<\mathfrak{bs}$, $\langle B\rangle\cap Z_{\alpha}\in\mathcal{I}^{+}(\mathcal{A})$ or $\langle B\rangle\cap Z_{\alpha}^{c}\in\mathcal{I}^{+}(\mathcal{A})$.
  2. (ii)
    If $\alpha<\mathfrak{bs}$ and $\langle B\rangle\cap Z_{\alpha}\in\mathcal{I}^{+}(\mathcal{A})$, then there exists some $C_{\alpha}^{0}\in\mathcal{H}^{-}(\mathcal{A})$ such that $\langle C_{\alpha}^{0}\rangle\subseteq\langle B\rangle\cap Z_{\alpha}$.
  3. (iii)
    If $\alpha<\mathfrak{bs}$ and $\langle B\rangle\cap Z_{\alpha}^{c}\in\mathcal{I}^{+}(\mathcal{A})$, then there exists some $C_{\alpha}^{1}\in\mathcal{H}^{-}(\mathcal{A})$ such that $\langle C_{\alpha}^{1}\rangle\subseteq\langle B\rangle\cap Z_{\alpha}^{c}$.
  4. (iv)
    There exists some $\alpha<\mathfrak{bs}$ such that $\langle B\rangle\cap Z_{\alpha}\in\mathcal{I}^{+}(\mathcal{A})$ and $\langle B\rangle\cap Z_{\alpha}^{c}\in\mathcal{I}^{+}(\mathcal{A})$.
Proof.
Let $\{S_{\alpha}:\alpha<\mathfrak{bs}\}\subseteq[\omega]^{\omega}$ be a block-splitting family, and let $Z_{\alpha}:=\langle(e_{n})_{n\in S_{\alpha}}\rangle$. We remark that $\langle(e_{n})_{n\notin S_{\alpha}}\rangle\subseteq Z_{\alpha}^{c}$. We shall show that $\{Z_{\alpha}:\alpha<\mathfrak{bs}\}$ satisfies the four properties above.
Property (i): Suppose that for all $\leq B, CC$ is compatible with some element of $\mathcal{A}$. Suppose that $\langle B\rangle\cap Z_{\alpha}\in\mathcal{I}(\mathcal{A})$ and $\langle B\rangle\cap Z_{\alpha}^{c}\in\mathcal{I}(\mathcal{A})$. Let $\{A_{i}:i<n\}\subseteq\mathcal{A}$ such that $(\langle B\rangle\cap Z_{\alpha})\setminus\bigcup_{i<n}\langle A_{i}\rangle$ and $(\langle B\rangle\cap Z_{\alpha}^{c})\setminus\bigcup_{i<n}\langle A_{i}\rangle$ are small. Since $B\in\mathcal{H}^{-}(\mathcal{A})$, there exists some $C\leq B$ such that:
$\langle C\rangle\subseteq\langle B\rangle\setminus\bigcup_{i<n}\langle A_{i}\rangle.$
However, since:
$\langle C\rangle\cap Z_{\alpha}\subseteq(\langle B\rangle\cap Z_{\alpha})\setminus\bigcup_{i<n}\langle A_{i}\rangle,$
$\langle C\rangle\cap Z_{\alpha}^{c}\subseteq(\langle B\rangle\cap Z_{\alpha}^{c})\setminus\bigcup_{i<n}\langle A_{i}\rangle,$
we have that $\langle C\rangle\cap Z_{\alpha}$ and $\langle C\rangle\cap Z_{\alpha}^{c}$ are small. This contradicts Lemma 5.4.
Property (ii): This follows from Lemma 2.3, where we may let $C_{0}^{\alpha}\in E^{[\infty]}$ be such that $\langle C_{0}^{\alpha}\rangle=\langle B\rangle\cap\langle(e_{n})_{n\in S_{\alpha}}\rangle$.
Property (iii): Suppose that $\langle B\rangle\cap Z_{\alpha}^{c}\in\mathcal{I}^{+}(\mathcal{A})$. If there exists some $C\in\mathcal{R}$ such that $\langle C\rangle\subseteq\langle B\rangle\cap Z_{\alpha}^{c}$ which is almost disjoint from every element of $\mathcal{A}$, then $C\in\mathcal{H}^{-}(\mathcal{A})$ and we are done, so assume otherwise. In a way similar to the one presented in the proof of Proposition 4.12, we define $A_{n},B_{n}$ for $n<\omega$ such that $\bigcup_{n<\omega}\langle B_{n}\rangle\subseteq\langle B\rangle\cap Z_{\alpha}^{c}$, $\{B_{n}:n<\omega\}$ is a disjoint family, and $A_{n}$ is compatible with $B_{n}$ for all $n$. Let $C_{n}$ be such that $\langle C_{n}\rangle\subseteq\langle A_{n}\rangle\cap\langle B_{n}\rangle$.
For each $n$, we write $C_{n}:=(x_{m}^{n})_{m<\omega}$. Since $\langle C_{n}\rangle\subseteq Z_{\alpha}^{c}$, this means that for all $m$, there exists some $l_{m,n}\notin S_{\alpha}$ and some $y_{m}^{n},z_{m}^{n}\in\langle C_{n}\rangle\cup\{0\}$ such that $y_{m}^{n}<e_{l_{m,n}}<z_{m}^{n}$, and $x_{m}^{n}=y_{m}^{n}+\lambda_{m,n}e_{l_{m,n}}+z_{m}^{n}$ for some $\lambda_{m,n}\neq 0$. Now $\Gamma:\omega\to\omega\times\omega$ be any function such that:
  1. (1)
    for all $n$, there are infinitely many $k$ such that $\Gamma(k)=(n,m)$ for some $m$, and;
  2. (2)
    if $k<k^{\prime}$, $\Gamma(k)=(n,m)$ and $\Gamma(k^{\prime})=(n^{\prime},m^{\prime})$, then $x_{n}^{m}<x_{n^{\prime}}^{m^{\prime}}$.
This implies that $C:=\left(x_{\Gamma(k)(1)}^{\Gamma(k)(0)}\right)_{k<\omega}$ is a well-defined infinite block sequence, where $\Gamma(k)=(\Gamma(k)(0),\Gamma(k)(1))$ such that $C$ is compatible with every $C_{n}$, so $C\in\mathcal{H}^{-}(\mathcal{A})$. Furthermore, we have that:
$\langle C\rangle\subseteq\left\langle\bigcup_{n<\omega}\langle C_{n}\rangle\right\rangle\subseteq\langle B\rangle,$
so $C\leq B$. Finally, if $x\in\langle C\rangle$, then $x=y+\lambda e_{l}+z$ for some $\lambda\neq 0$, so $x\notin Z_{\alpha}$. In other words, $\langle C\rangle\subseteq Z_{\alpha}^{c}$. Therefore, $C\in\mathcal{H}^{-}(\mathcal{A})$ and $\langle C\rangle\subseteq\langle B\rangle\cap Z_{\alpha}^{c}$, as desired.
Property (iv): We first suppose that $B$ is almost disjoint from every element of $\mathcal{A}$. If $B=(x_{n})_{n<\omega}$, we define, for $n<\omega$:
$P_{0}$ $:=\{0,\dots,\max(\operatorname{supp}(x_{n}))\},$
$P_{n+1}$ $:=\{\max(\operatorname{supp}(x_{n}))+1,\dots,\max(\operatorname{supp}(x_{n+1})\}.$
We have that $\mathcal{P}:=\{P_{n}:n<\omega\}$ is an interval partition such that $\operatorname{supp}(x_{n})\subseteq P_{n}$ for all $n$. Since $\{S_{\alpha}:\alpha<\mathfrak{bs}\}$ is a block-splitting family, there exists some $\alpha<\mathfrak{bs}$ that splits $\mathcal{P}$. Let $T_{0}:=\{n<\omega:P_{n}\subseteq S_{\alpha}\}$ and $T_{1}:=\{n<\omega:P_{n}\cap S_{\alpha}^{c}\}$. Define $B_{0}:=(x_{n})_{n\in T_{0}}$ and $B_{1}:=(x_{n})_{n\in T_{1}}$. Since $B_{0},B_{1}\leq B$, and $B$ is almost disjoint from every element of $\mathcal{A}$, so are $B_{0},B_{1}$ and so $B_{0},B_{1}\in\mathcal{H}^{-}(\mathcal{A})$. If $n\in T_{0}$, then $\operatorname{supp}(x_{n})\subseteq P_{n}\subseteq S_{\alpha}$, $x_{n}\in(e_{n})_{n\in S_{\alpha}}$. Therefore, $\langle B_{0}\rangle\subseteq\langle B\rangle\cap Z_{\alpha}$. By a similar reasoning, we may also conclude that $\langle B_{1}\rangle\subseteq\langle B\rangle\cap Z_{\alpha}^{c}$.
Now suppose that there is no $C\leq B$ which is almost disjoint from every element of $\mathcal{A}$ (if such a $C$ exists, we may apply the argument in the previous paragraph to $C$, and use the fact that $\mathcal{I}^{+}(\mathcal{A})$ is $\subseteq$-upward closed). By Lemma 4.11, let $\{A_{n}:n<\omega\}\subseteq\mathcal{A}$ be an infinite set and let $\{B_{n}:n<\omega\}$ be an infinite disjoint family such that $\langle B_{n}\rangle\subseteq\langle B\rangle\cap\langle A_{n}\rangle$ for all $n$. Let $\Gamma:\omega\to\omega$ be an infinite-to-one surjection. We define an infinite block sequence $(x_{n})_{n<\omega}$ inductively as follows: If $(x_{k})_{k<n}$ has been defined, let $x_{n}\in\langle B_{\Gamma(n)}\rangle$ be any vector $(x_{k})_{k<n}<x_{n}$. Having defined $(x_{n})_{n<\omega}$, we now define an interval partition $\mathcal{P}:=\{P_{m}:m<\omega\}$ as follows: Let $P_{0}:=\{0,\dots,\max(\operatorname{supp}(x_{n_{0}}))\}$, where $n_{0}$ is the least integer such that $\Gamma(n_{0})=0$. If $P_{m}$ has been defined with $\max(P_{m})=n_{m}$, Let $\min(P_{m+1})=n_{m}+1$, and let $\max(P_{m+1})$ be large enough so that for all $i\leq m+1$, there exists some $n$ such that $\Gamma(n)=i$ and $\operatorname{supp}(x_{n})\subseteq P_{m+1}$. This is possible as $\Gamma$ is infinite-to-one.
Now let $\alpha<\mathfrak{bs}$ be such that $S_{\alpha}$ splits $\mathcal{P}$. For each $i$, let $T_{i}^{0}:=\{n<\omega:\Gamma(n)=i\wedge\operatorname{supp}(x_{n})\subseteq S_{\alpha}\}$ and $T_{i}^{1}:=\{n<\omega:\Gamma(n)=i\wedge\operatorname{supp}(x_{n})\cap S_{\alpha}=\emptyset\}$. By our construction of $\mathcal{P}$, we have that $T_{i}^{0}$ and $T_{i}^{1}$ are infinite for all $i$. By a similar reasoning to the first case, we may conclude that $\langle B\rangle\cap Z_{\alpha}$ and $\langle B\rangle\cap Z_{\alpha}^{c}$ are compatible with $A_{i}$ for all $i$, so they are both in $\mathcal{I}^{+}(\mathcal{A})$. ∎
We are now ready to prove the theorem.
Proof of Theorem 1.5.
Let $\{Z_{\alpha}:\alpha<\mathfrak{bs}\}$ be a family of subsets of vectors satisfying the conclusions of Lemma 5.5. We write $Z_{\alpha}^{0}:=Z_{\alpha}$ and $Z_{\alpha}^{1}:=Z_{\alpha}^{c}$ for all $\alpha$. With this, given any almost disjoint family $\mathcal{A}\subseteq E^{[\infty]}$ and $B\in\mathcal{H}^{-}(\mathcal{A})$, there exist $\alpha<\mathfrak{bs}$ and $\tau_{B}^{\mathcal{A}}\in 2^{\alpha}$ such that:
  1. (1)
    If $\beta<\alpha$, then $\langle B\rangle\cap Z_{\beta}^{1-\tau_{B}^{\mathcal{A}}(\beta)}\in\mathcal{I}(\mathcal{A})$ (so $\langle B\rangle\cap Z_{\beta}^{\tau_{B}^{\mathcal{A}}(\beta)}\in\mathcal{I}^{+}(\mathcal{A})$ by property (i) of Lemma 5.5);
  2. (2)
    $\langle B\rangle\cap Z_{\alpha}^{0}\in\mathcal{I}^{+}(\mathcal{A})$ and $\langle B\rangle\cap Z_{\alpha}^{1}\in\mathcal{I}^{+}(\mathcal{A})$ (such an $\alpha$ exists by property (iv) of Lemma 5.5).
Observe that for a fixed almost disjoint family $\mathcal{A}$, $\tau_{B}^{\mathcal{A}}\in 2^{<\mathfrak{bs}}$ is unique, and if $B^{\prime}\leq B$ and $B^{\prime}\in\mathcal{H}^{-}(\mathcal{A})$, then $\tau_{B}^{\mathcal{A}}\subseteq\tau_{B^{\prime}}^{\mathcal{A}}$. Let $\operatorname{Lim}(\mathfrak{c})$ be the set of limit ordinals below $\mathfrak{c}$. Let $\{Y_{\alpha}:\alpha<\operatorname{Lim}(\mathfrak{c})\}$ be an enumeration of $\mathcal{P}(E)$, and let $\{B_{\alpha}:\alpha<\operatorname{Lim}(\mathfrak{c})\}$ be an enumeration of $E^{[\infty]}$ in a way that for any $B\in E^{[\infty]}$ and $Y\subseteq E$, there are cofinally many $\alpha<\operatorname{Lim}(\mathfrak{c})$ such that $(Y_{\alpha},B_{\alpha})=(Y,B)$. We shall define our mad family $\mathcal{A}=\{A_{\alpha}:\alpha<\mathfrak{c}\}$ recursively, where $\mathcal{A}_{\alpha}:=\{A_{\beta}:\beta<\alpha\}$, along with a collection $\{A_{\alpha}^{\prime}:\alpha<\operatorname{Lim}(\mathfrak{c})\}$ of block sequences, and an injective sequence $\{\sigma_{\alpha}:\alpha<\operatorname{Lim}(\mathfrak{c})\}$ of elements of $2^{<\mathfrak{bs}}$ such that the following holds:
  1. (1)
    If $\beta<\alpha$ is a limit ordinal, then $\sigma_{\alpha}\not\sqsubseteq\sigma_{\beta}$.
  2. (2)
    $A_{\alpha}^{\prime}$ is almost disjoint with $A_{\beta}$ for all $\beta<\alpha$.
  3. (3)
    If $\xi<\operatorname{dom}(\sigma_{\alpha})$, then $\langle A_{\alpha}^{\prime}\rangle\setminus Z_{\xi}^{\sigma_{\alpha}(\xi)}$ is small (hence, by Lemma 5.4, is very small).
  4. (4)
    $\{A_{\alpha+n}:n<\omega\}$ is a disjoint family below $A_{\alpha}^{\prime}$.
For our base case, we let $\mathcal{A}_{\omega}$ be any countable almost disjoint family. Assume that $\alpha\geq\omega$ is a limit ordinal, and that $A_{\beta}$, $A_{\beta}^{\prime}$ and $\sigma_{\beta}$ have been defined for $\beta<\alpha$, and satisfy the requirements stated above so far. We define the element $B:=B_{\alpha}$ depending on the following scenarios:
  1. (a)
    If $B_{\alpha}\notin\mathcal{H}(\mathcal{A}_{\alpha})$, then let $B:=B_{\alpha}$ if $B_{\alpha}\in\mathcal{H}^{-}(\mathcal{A}_{\alpha})$. Otherwise, let $B$ be any element of $\mathcal{H}^{-}(\mathcal{A}_{\alpha})$.
  2. (b)
    Suppose that $B_{\alpha}\in\mathcal{H}(\mathcal{A}_{\alpha})$, and:
    $\mathcal{C}_{\alpha}:=\{A\in\mathcal{A}_{\alpha}:B_{\alpha}\text{ and }A\text{ are compatible}\}.$
    Suppose further that there exists some $B_{\alpha}^{\prime}\leq B_{\alpha}$ that is almost disjoint from every element of $\mathcal{C}_{\alpha}$, and $\langle B_{\alpha}^{\prime}\rangle\subseteq Y_{\alpha}$. Then we let $B:=B_{\alpha}^{\prime}$.
  3. (c)
    Suppose that $B_{\alpha}\in\mathcal{H}(\mathcal{A}_{\alpha})$, and for all $B\leq B_{\alpha}$ which is almost disjoint from every element of $\mathcal{C}_{\alpha}$, there is no $B^{\prime}\leq B$ such that $\langle B^{\prime}\rangle\not\subseteq Y_{\alpha}$. Then we let $B:=B_{\alpha}$.
We inductively define $\{C_{s}:s\in 2^{<\omega}\}\subseteq\mathcal{H}^{-}(\mathcal{A}_{\alpha})$, $\{\eta_{s}:s\in 2^{<\omega}\}\subseteq 2^{<\mathfrak{bs}}$ and ordinals $\{\alpha_{s}:s\in 2^{<\omega}\}$ as follows:
  1. (1)
    $C_{\emptyset}:=B$.
  2. (2)
    $\eta_{s}:=\tau_{C_{s}}^{\mathcal{A}_{\alpha}}$ and $\alpha_{s}:=\operatorname{dom}(\eta_{s})$.
  3. (3)
    $C_{s^{\frown}0},C_{s^{\frown}1}\in\mathcal{H}^{-}(\mathcal{A})$ are such that $\langle C_{s^{\frown}0}\rangle\subseteq\langle C_{s}\rangle\cap Z_{\alpha_{s}}^{0}$ and $\langle C_{s^{\frown}1}\rangle\subseteq\langle C_{s}\rangle\cap Z_{\alpha_{s}}^{1}$ (which both exist by properties (ii) and (iii) of Lemma 5.5).
We then define $\eta_{f}:=\bigcup_{n<\omega}\eta_{f\mathord{\upharpoonright}n}$ for any $f\in 2^{\omega}$, and note that $\eta_{f}\in 2^{<\mathfrak{bs}}$ as $\mathfrak{bs}$ has uncountable cofinality. We note the following observations:
  1. (1)
    If $s\sqsubseteq t$, then $C_{s}\leq C_{t}$ and $\eta_{s}\sqsubseteq\eta_{t}$.
  2. (2)
    If $s$ and $t$ are incompatible, then so are $\eta_{s}$ and $\eta_{t}$.
  3. (3)
    Consequently, if $f$ and $g$ are incompatible, then so are $\eta_{f}$ and $\eta_{g}$.
Since $\alpha<\mathfrak{c}$, we may let $f\in 2^{\omega}$ be such that $\eta_{f}\not\sqsubseteq\sigma_{\beta}$ for all limit ordinals $\beta<\alpha$. We now have a $\leq$-decreasing sequence $(C_{f\mathord{\upharpoonright}n})_{n<\omega}$ in $\mathcal{H}^{-}(\mathcal{A}_{\alpha})$, so by Proposition 4.12 there exists some $D\in\mathcal{H}^{-}(\mathcal{A}_{\alpha})$ such that $D\leq^{*}C_{f\mathord{\upharpoonright}n}$ for all $n$.
Let $\sigma_{\alpha}:=\eta_{f}$. Observe that for all $\xi<\operatorname{dom}(\eta_{f})$, if $m<\omega$ is such that $\xi\in\operatorname{dom}(\eta_{f\mathord{\upharpoonright}n})$, and $D/N\leq C_{f\mathord{\upharpoonright}n}$, then:
$\langle D/N\rangle\cap Z_{\xi}^{1-\eta_{f}(\xi)}=\langle D/N\rangle\cap Z_{\xi}^{1-\eta_{f\mathord{\upharpoonright}n}(\xi)}\subseteq\langle C_{f\mathord{\upharpoonright}n}\rangle\cap Z_{\xi}^{1-\eta_{f\mathord{\upharpoonright}n}(\xi)}\in\mathcal{I}(\mathcal{A}_{\alpha}),$
so that for all $\xi<\operatorname{dom}(\sigma_{\alpha})$, $\langle D\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)}\in\mathcal{I}(\mathcal{A})$ (as $\langle D\rangle\setminus\langle D/N\rangle$ is very small, by Lemma 5.4). Therefore, for each $\xi<\operatorname{dom}(\sigma_{\alpha})$ we may find a finite subset $\mathcal{F}_{\xi}\subseteq\mathcal{A}_{\alpha}$ such that $\left(\langle D\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)}\right)\setminus\bigcup_{A\in\mathcal{F}_{\xi}}\langle A\rangle$ is small. Now let:
$\mathcal{B}_{0}$ $:=\{A_{\beta+n}:\sigma_{\beta}\sqsubseteq\sigma_{\alpha}\},$
$\mathcal{B}_{1}$ $:=\bigcup_{\xi<\operatorname{dom}(\sigma_{\alpha})}\mathcal{F}_{\xi},$
$\mathcal{B}$ $:=\mathcal{B}_{0}\cup\mathcal{B}_{1}.$
By Lemma 2.3, $\{\langle B^{\prime}\rangle\cap\langle D\rangle:B^{\prime}\in\mathcal{B}\}$ is an almost disjoint family of block subspaces of $\langle D\rangle$. By Lemma 5.3, $|\mathcal{B}|<\mathfrak{bs}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$, so there exists some $A_{\alpha}^{\prime}\leq D$ such that $A_{\alpha}^{\prime}$ is almost disjoint from every element of $\mathcal{B}$. Note that $A_{\alpha}^{\prime}$ may be chosen to be a block subspace as every infinite-dimensional subspace contains a block subspace.
Claim 8.
  1. (1)
    $\langle A_{\alpha}^{\prime}\rangle\setminus Z_{\xi}^{\sigma_{\alpha}(\xi)}$ is small.
  2. (2)
    For all $\beta<\alpha$, $A_{\alpha}^{\prime}$ and $A_{\beta}$ are almost disjoint.
Proof.
  1. (1)
    Note that $\langle A_{\alpha}^{\prime}\rangle\setminus Z_{\xi}^{\sigma_{\alpha}(\xi)}=\langle A_{\alpha}^{\prime}\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)}$. We have that $\langle A_{\alpha}^{\prime}\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)}\setminus\bigcup_{A\in\mathcal{F}_{\xi}}\langle A\rangle$ is small by above, and since $A_{\alpha}^{\prime}$ is almost disjoint from every $A\in\mathcal{F}_{\xi}$, we have that:
    $\langle A_{\alpha}^{\prime}\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)}\cap\bigcup_{A\in\mathcal{F}_{\xi}}\langle A\rangle$ $\subseteq\bigcup_{A\in\mathcal{F}_{\xi}}\langle A_{\alpha}^{\prime}\rangle\cap\langle A\rangle.$
    By Lemma 5.4, $\langle A_{\alpha}\rangle\cap\langle A\rangle$ is very small for each $A\in\mathcal{F}_{\xi}$, so $\langle A_{\alpha}^{\prime}\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)}$ is small.
  2. (2)
    Let $\beta+n<\alpha$, where $\beta$ is a limit ordinal. If $A_{\beta+n}\in\mathcal{B}_{0}$, then $A_{\alpha}$ is almost disjoint from $A_{\beta+n}$ by construction, so we are done. Now assume that $A_{\beta+n}\notin W$. Let $\xi$ be the least ordinal such that $\sigma_{\beta}(\xi)\neq\sigma_{\alpha}(\xi)$. By induction hypothesis, we have that $W:=\langle A_{\beta}^{\prime}\rangle\setminus Z_{\xi}^{\sigma_{\beta}(\xi)}=\langle A_{\beta}^{\prime}\rangle\setminus Z_{\xi}^{1-\sigma_{\alpha}(\xi)}$ is small (and hence very small). Since $A_{\beta+n}\leq A_{\beta}^{\prime}$, we have that for some small set $W^{\prime}$:
    $\langle A_{\alpha}^{\prime}\rangle\cap\langle A_{\beta+n}\rangle$ $\subseteq W\cup(\langle A_{\alpha}^{\prime}\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)})$
    $\subseteq W\cup W^{\prime}\cup\bigcup_{A\in F_{\xi}}\langle A_{\alpha}^{\prime}\rangle\cap\langle A\rangle,$
    which is also small as all the components in the union are very small, as desired.
We complete the induction by letting $\{A_{\alpha+n}:n<\omega\}$ be any infinite disjoint family below $A_{\alpha}^{\prime}$. We note that $\mathcal{A}$ is a mad family – otherwise, if $B$ is almost disjoint from every element of $\mathcal{A}$, then $B\in\mathcal{H}^{-}(\mathcal{A}_{\alpha})$ for all $\alpha<\mathfrak{c}$. Let $\alpha$ be any limit ordinal such that $B_{\alpha}=B$, and we have that $A_{\alpha}\leq B$, a contradiction. It remains to show that $\mathcal{A}$ is full.
Let $Y\subseteq E$, and suppose that $Y$ is $\mathcal{H}(\mathcal{A})$-dense below some $B\in\mathcal{H}(\mathcal{A})$. Since $\mathcal{H}(\mathcal{A})=\bigcup_{\alpha<\mathfrak{c}}\mathcal{H}(\mathcal{A}_{\alpha})$, we may let $\alpha$ be large enough so that $(Y_{\alpha},B_{\alpha})=(Y,B)$ and $B\in\mathcal{H}(\mathcal{A}_{\alpha})$. This means that in the induction step, we have either case (b) or (c). Observe that (c) is not possible – $A_{\alpha}^{\prime}\leq B_{\alpha}$ is almost disjoint from every element of $\mathcal{C}_{\alpha}$, and $A_{\alpha}^{\prime}\in\mathcal{H}(\mathcal{A})$ as $A_{\alpha}$ is compatible with $A_{\alpha+n}$ for all $n$. However, there is no $B^{\prime}\leq A_{\alpha}^{\prime}$ in which $\langle B^{\prime}\rangle\subseteq Y_{\alpha}$, contradicting that $Y$ is $\mathcal{H}(\mathcal{A})$-dense below $B$. This means that case (b) must have occurred in this induction step. Since $B_{\alpha^{\prime}}$ is compatible with $A_{\alpha+n}$ for all $n$, $B_{\alpha^{\prime}}\in\mathcal{H}(\mathcal{A})$ and $\langle B_{\alpha}^{\prime}\rangle\subseteq\langle Y_{\alpha}\rangle$, as desired. ∎
5.3. Completely separable mad family of block subspaces
We dedicate this section to proving Theorem 1.6.
Definition 5.6.
A mad family $\mathcal{A}\subseteq E^{[\infty]}$ is completely separable if for all $Y\subseteq E$ such that $\langle Y\rangle\in\mathcal{I}^{++}(\mathcal{A})$, there exists some $A\in\mathcal{A}$ such that $\langle A\rangle\subseteq\langle Y\rangle$.
The proof of Theorem 1.5 may be adjusted to obtain a completely separable mad family of block subspaces, with the main change being at the induction step. We describe the details here.
Proof of Theorem 1.6.
Let $\{Z_{\alpha}:\alpha<\mathfrak{bs}\}$ be a family of subsets of vectors satisfying the conclusions of Lemma 5.5. We write $Z_{\alpha}^{0}:=Z_{\alpha}$ and $Z_{\alpha}^{1}:=Z_{\alpha}^{c}$ for all $\alpha$. With this, given any almost disjoint family $\mathcal{A}\subseteq E^{[\infty]}$ and $B\in\mathcal{H}^{-}(\mathcal{A})$, there exist $\alpha<\mathfrak{bs}$ and $\tau_{B}^{\mathcal{A}}\in 2^{\alpha}$ such that:
  1. (1)
    If $\beta<\alpha$, then $\langle B\rangle\cap Z_{\beta}^{1-\tau_{B}^{\mathcal{A}}(\beta)}\in\mathcal{I}(\mathcal{A})$ (so $\langle B\rangle\cap Z_{\beta}^{\tau_{B}^{\mathcal{A}}(\beta)}\in\mathcal{I}^{+}(\mathcal{A})$ by property (i) of Lemma 5.5);
  2. (2)
    $\langle B\rangle\cap Z_{\alpha}^{0}\in\mathcal{I}^{+}(\mathcal{A})$ and $\langle B\rangle\cap Z_{\alpha}^{1}\in\mathcal{I}^{+}(\mathcal{A})$ (such an $\alpha$ exists by property (iv) of Lemma 5.5).
Observe that for a fixed almost disjoint family $\mathcal{A}$, $\tau_{B}^{\mathcal{A}}\in 2^{<\mathfrak{bs}}$ is unique, and if $B^{\prime}\leq B$ and $B^{\prime}\in\mathcal{H}^{-}(\mathcal{A})$, then $\tau_{B}^{\mathcal{A}}\subseteq\tau_{B^{\prime}}^{\mathcal{A}}$ We may need that for any $Y\subseteq E$, there are cofinally many $\alpha<\mathfrak{c}$ such that $Y_{\alpha}=Y$. We shall define our mad family $\mathcal{A}=\{A_{\alpha}:\alpha<\mathfrak{c}\}$ recursively, where $\mathcal{A}_{\alpha}:=\{A_{\beta}:\beta<\alpha\}$, along with an injective sequence $\{\sigma_{\alpha}:\alpha<\mathfrak{c}\}$ of elements of $2^{<\mathfrak{bs}}$ such that the following holds:
  1. (1)
    If $\beta<\alpha$, then $\sigma_{\alpha}\not\sqsubseteq\sigma_{\beta}$.
  2. (2)
    If $\xi<\operatorname{dom}(\sigma_{\alpha})$, then $\langle A_{\alpha}\rangle\setminus Z_{\xi}^{\sigma_{\alpha}(\xi)}$ is small (hence, by Lemma 5.4, is very small).
For our base case, we let $\mathcal{A}_{\omega}$ be any countable almost disjoint family. Assume that $A_{\beta}$, $\sigma_{\beta}$ have been defined for $\beta<\alpha$ such that they satisfy the requirements stated above so far. We inductively define $\{C_{s}:s\in 2^{<\omega}\}\subseteq\mathcal{H}^{-}(\mathcal{A}_{\alpha})$, $\{\eta_{s}:s\in 2^{<\omega}\}\subseteq 2^{<\mathfrak{bs}}$ and ordinals $\{\alpha_{s}:s\in 2^{<\omega}\}$ as follows:
  1. (1)
    If $Y_{\alpha}\in\mathcal{I}^{+}(\mathcal{A})$, let $C_{\emptyset}\in\mathcal{H}^{-}(\mathcal{A})$ such that $\langle C_{\emptyset}\rangle\subseteq Y_{\alpha}$. This is possible by Lemma 5.5. If $Y_{\alpha}\notin\mathcal{I}^{+}(\mathcal{A})$, let $C_{\emptyset}\in\mathcal{H}^{-}(\mathcal{A})$ be any bloc subspace.
  2. (2)
    $\eta_{s}:=\tau_{C_{s}}^{\mathcal{A}_{\alpha}}$ and $\alpha_{s}:=\operatorname{dom}(\eta_{s})$.
  3. (3)
    $C_{s^{\frown}0},C_{s^{\frown}1}\in\mathcal{H}^{-}(\mathcal{A})$ are such that $\langle C_{s^{\frown}0}\rangle\subseteq\langle C_{s}\rangle\cap Z_{\alpha_{s}}^{0}$ and $\langle C_{s^{\frown}1}\rangle\subseteq\langle C_{s}\rangle\cap Z_{\alpha_{s}}^{1}$ (which both exist by properties (ii) and (iii) of Lemma 5.5).
We then define $\eta_{f}:=\bigcup_{n<\omega}\eta_{f\mathord{\upharpoonright}n}$ for any $f\in 2^{\omega}$, and note that $\eta_{f}\in 2^{<\mathfrak{bs}}$ as $\mathfrak{bs}$ has uncountable cofinality. We note the following observations:
  1. (1)
    If $s\sqsubseteq t$, then $C_{s}\leq C_{t}$ and $\eta_{s}\sqsubseteq\eta_{t}$.
  2. (2)
    If $s$ and $t$ are incompatible, then so are $\eta_{s}$ and $\eta_{t}$.
  3. (3)
    Consequently, if $f$ and $g$ are incompatible, then so are $\eta_{f}$ and $\eta_{g}$.
Since $\alpha<\mathfrak{c}$, we may let $f\in 2^{\omega}$ be such that $\eta_{f}\not\sqsubseteq\sigma_{\beta}$ for all limit ordinals $\beta<\alpha$. We now have a $\leq$-decreasing sequence $(C_{f\mathord{\upharpoonright}n})_{n<\omega}$ in $\mathcal{H}^{-}(\mathcal{A}_{\alpha})$, so by Proposition 4.12 there exists some $D\in\mathcal{H}^{-}(\mathcal{A}_{\alpha})$ such that $D\leq^{*}C_{f\mathord{\upharpoonright}n}$ for all $n$.
Let $\sigma_{\alpha}:=\eta_{f}$. Observe that for all $\xi<\operatorname{dom}(\eta_{f})$, if $m<\omega$ is such that $\xi\in\operatorname{dom}(\eta_{f\mathord{\upharpoonright}n})$, and $D/N\leq C_{f\mathord{\upharpoonright}n}$, then:
$\langle D/N\rangle\cap Z_{\xi}^{1-\eta_{f}(\xi)}=\langle D/N\rangle\cap Z_{\xi}^{1-\eta_{f\mathord{\upharpoonright}n}(\xi)}\subseteq\langle C_{f\mathord{\upharpoonright}n}\rangle\cap Z_{\xi}^{1-\eta_{f\mathord{\upharpoonright}n}(\xi)}\in\mathcal{I}(\mathcal{A}_{\alpha}),$
so that for all $\xi<\operatorname{dom}(\sigma_{\alpha})$, $\langle D\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)}\in\mathcal{I}(\mathcal{A})$ (as $\langle D\rangle\setminus\langle D/N\rangle$ is very small, by Lemma 5.4). Therefore, for each $\xi<\operatorname{dom}(\sigma_{\alpha})$ we may find a finite subset $\mathcal{F}_{\xi}\subseteq\mathcal{A}_{\alpha}$ such that $\left(\langle D\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)}\right)\setminus\bigcup_{A\in\mathcal{F}_{\xi}}\langle A\rangle$ is small. Now let:
$\mathcal{B}_{0}$ $:=\{A_{\beta+n}:\sigma_{\beta}\sqsubseteq\sigma_{\alpha}\},$
$\mathcal{B}_{1}$ $:=\bigcup_{\xi<\operatorname{dom}(\sigma_{\alpha})}\mathcal{F}_{\xi},$
$\mathcal{B}$ $:=\mathcal{B}_{0}\cup\mathcal{B}_{1}.$
By Lemma 2.3, $\{\langle B^{\prime}\rangle\cap\langle D\rangle:B^{\prime}\in\mathcal{B}\}$ is an almost disjoint family of block subspaces of $\langle D\rangle$. By Lemma 5.3, $|\mathcal{B}|<\mathfrak{bs}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$, so there exists some $A_{\alpha}^{\prime}\leq D$ such that $A_{\alpha}^{\prime}$ is almost disjoint from every element of $\mathcal{B}$. Note that $A_{\alpha}^{\prime}$ may be chosen to be a block subspace as every infinite-dimensional subspace contains a block subspace.
Claim 9.
  1. (1)
    $\langle A_{\alpha}\rangle\setminus Z_{\xi}^{\sigma_{\alpha}(\xi)}$ is small.
  2. (2)
    For all $\beta<\alpha$, $A_{\alpha}$ and $A_{\beta}$ are almost disjoint.
Proof.
  1. (1)
    Note that $\langle A_{\alpha}\rangle\setminus Z_{\xi}^{\sigma_{\alpha}(\xi)}=\langle A_{\alpha}\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)}$. We have that $\langle A_{\alpha}\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)}\setminus\bigcup_{A\in\mathcal{F}_{\xi}}\langle A\rangle$ is small by above, and since $A_{\alpha}$ is almost disjoint from every $A\in\mathcal{F}_{\xi}$, we have that:
    $\langle A_{\alpha}\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)}\cap\bigcup_{A\in\mathcal{F}_{\xi}}\langle A\rangle$ $\subseteq\bigcup_{A\in\mathcal{F}_{\xi}}\langle A_{\alpha}\rangle\cap\langle A\rangle.$
    By Lemma 5.4, $\langle A_{\alpha}\rangle\cap\langle A\rangle$ is very small for each $A\in\mathcal{F}_{\xi}$, so $\langle A_{\alpha}\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)}$ is small.
  2. (2)
    Let $\beta<\alpha$. If $A_{\beta}\in\mathcal{B}_{0}$, then $A_{\alpha}$ is almost disjoint from $A_{\beta}$ by construction, so we are done. Now assume that $A_{\beta}\notin W$. Let $\xi$ be the least ordinal such that $\sigma_{\beta}(\xi)\neq\sigma_{\alpha}(\xi)$. By induction hypothesis, we have that $W:=\langle A_{\beta}\rangle\setminus Z_{\xi}^{\sigma_{\beta}(\xi)}=\langle A_{\beta}\rangle\setminus Z_{\xi}^{1-\sigma_{\alpha}(\xi)}$ is small (and hence very small). Thus, we have that for some small set $W^{\prime}$:
    $\langle A_{\alpha}\rangle\cap\langle A_{\beta}\rangle$ $\subseteq W\cup(\langle A_{\alpha}\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)})$
    $\subseteq W\cup W^{\prime}\cup\bigcup_{A\in F_{\xi}}\langle A_{\alpha}\rangle\cap\langle A\rangle,$
    which is also small as all the components in the union are very small, as desired.
This completes the induction. If $\langle Y\rangle\in\mathcal{I}^{+}(\mathcal{A})$, then $\mathcal{I}^{+}(\mathcal{A}_{\alpha})$ for all $\alpha$, and there exists some $\alpha$ such that $Y_{\alpha}=\langle Y\rangle$. This implies that $A_{\alpha}\subseteq\langle Y\rangle$, as desired. Setting $\langle Y\rangle=\langle B\rangle$ for some $B\in E^{[\infty]}$, this implies that $\mathcal{A}$ is mad. ∎
6. Topological Ramsey spaces and the abstract Mathias forcing
We dedicate this section to an overview of the axioms of topological Ramsey spaces, and the abstract Mathias forcing developed in [DMN15].
6.1. Topological Ramsey spaces
We recap the four axioms presented by Todorčević in [T10], which are sufficient conditions for a triple $(\mathcal{R},\leq,r)$ to be a topological Ramsey space. Here, $\mathcal{R}$ is a non-empty set, $\leq$ be a quasi-order on $\mathcal{R}$, and $r:\mathcal{R}\times\omega\to\mathcal{AR}$ is a function. We also define a sequence of maps $r_{n}:\mathcal{R}\to\mathcal{AR}$ by $r_{n}(A):=r(A,n)$ for all $A\in\mathcal{R}$. Let $\mathcal{AR}_{n}\subseteq\mathcal{AR}$ be the image of $r_{n}$ (i.e. $a\in\mathcal{AR}_{n}$ iff $a=r_{n}(A)$ for some $A\in\mathcal{R}$).
The four axioms are as follows:
  1. (A1)
    1. (1)
      $r_{0}(A)=\emptyset$ for all $A\in\mathcal{AR}$.
    2. (2)
      $A\neq B$ implies $r_{n}(A)\neq r_{n}(B)$ for some $n$.
    3. (3)
      $r_{n}(A)=r_{m}(B)$ implies $n=m$ and $r_{k}(A)=r_{k}(B)$ for all $k<n$.
    For each $a\in\mathcal{AR}$, let $\operatorname{lh}(a)$ denote the unique $n$ in which $a\in\mathcal{AR}_{n}$. By Axiom A1(3), this $n$ is well-defined.
  2. (A2)
    There is a quasi-ordering $\leq_{\mathrm{fin}}$ on $\mathcal{AR}$ such that:
    1. (1)
      $\{a\in\mathcal{AR}:a\leq_{\mathrm{fin}}b\}$ is finite for all $b\in\mathcal{AR}$.
    2. (2)
      $A\leq B$ iff $\forall n\exists m[r_{n}(A)\leq_{\mathrm{fin}}r_{m}(B)]$.
    3. (3)
      $\forall a,b\in\mathcal{AR}[a\sqsubseteq b\wedge b\leq_{\mathrm{fin}}c\to\exists d\sqsubseteq c[a\leq_{\mathrm{fin}}d]]$.
  3. (A3)
    We may define the Ellentuck neighbourhoods as follows: For any $A\in\mathcal{R}$, $a\in\mathcal{AR}$ and $n\in\mathbb{N}$, we let:
    $[a,A]$ $:=\{B\in\mathcal{R}:B\leq A\wedge\exists n[r_{n}(B)=a]\},$
    $[n,A]$ $:=[r_{n}(A),A].$
    Then the depth function defined by, for $B\in\mathcal{R}$ and $a\in\mathcal{AR}$:
    $0pt_{B}(a):=\begin{cases}\min\{n<\omega:a\leq_{\mathrm{fin}}r_{n}(B)\},&\text{if such $n$ exists},\\ \infty,&\text{otherwise},\end{cases}$
    satisfies the following:
    1. (1)
      If $0pt_{B}(a)<\infty$, then for all $A\in[0pt_{B}(a),B]$, $a,A]\neq\emptyset$.
    2. (2)
      If $A\leq B$ and $a,A]\neq\emptyset$, then there exists $A^{\prime}\in[0pt_{B}(a),B]$ such that $\emptyset\neq[a,A^{\prime}]\subseteq[a,A]$.
    For each $A\in\mathcal{R}$, we let:
    $\mathcal{AR}\mathord{\upharpoonright}A:=\{a\in\mathcal{AR}:\exists n[a\leq_{\mathrm{fin}}r_{n}(A)]\}.$
    If $a\in\mathcal{AR}\mathord{\upharpoonright}A$, we also define:
    $\mathcal{AR}\mathord{\upharpoonright}[a,A]$ $:=\{b\in\mathcal{AR}\mathord{\upharpoonright}A:a\sqsubseteq b\},$
    $r_{n}[a,A]$ $:=\{b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]:\operatorname{lh}(b)=n\}.$
  4. (A4)
    If $0pt_{B}(a)<\infty$ and if $\mathcal{O}\subseteq\mathcal{AR}_{\operatorname{lh}(a)+1}$, then there exists $A\in[0pt_{B}(a),B]$ such that $r_{\operatorname{lh}(a)+1}[a,A]\subseteq\mathcal{O}$ or $r_{\operatorname{lh}(a)+1}[a,A]\subseteq\mathcal{O}^{c}$.
Note that by axiom A1, we may identify each element $A\in\mathcal{R}$ as a sequence of elements of $\mathcal{AR}$, via the map $A\mapsto(r_{n}(A))_{n<\omega}$. Therefore, we may identify $\mathcal{R}$ as a subset of $\mathcal{AR}^{\mathbb{N}}$.
Definition 6.1.
A triple $(\mathcal{R},\leq,r)$ is said to be a closed triple if $\mathcal{R}$ is a metrically closed subset of $\mathcal{AR}^{\mathbb{N}}$.
A closed triple satisfying $(\mathcal{R},\leq,r)$ satisfying A1-A4 is a topological Ramsey space, as it satisfies the abstract Ellentuck theorem – that is, a subset $\mathcal{X}\subseteq\mathcal{R}$ is Ramsey iff it has the property of Baire relative to the Ellentuck topology. This is the abstract Ellentuck theorem by Todorčević – see Theorem 5.4 of [T10].
We list some commonly used terminologies and notations for topological Ramsey spaces.
  • If $A\leq B$, we say that $A$ is a reduction of $B$.
  • $\mathcal{AR}\mathord{\upharpoonright}A:=\{b\in\mathcal{AR}:\exists n[b\leq_{\mathrm{fin}}r_{n}(A)]\}$.
  • $\mathcal{AR}\mathord{\upharpoonright}[a,A]:=\{b\in\mathcal{AR}:a\sqsubseteq b\wedge b\in\mathcal{AR}\mathord{\upharpoonright}A\}$.
Mathias’ original proof of his celebrated theorem (i.e. Theorem 1.7 for $\omega]^{\omega}$) uses the (original) Mathias forcing localised to a coideal $\mathcal{H}$. As we shall use a similar approach to prove Theorem 1.7, it is necessary to extend the notion of a coideal to topological Ramsey spaces.
Definition 6.2 (Definition 3.1, [DMN15]).
Let $(\mathcal{R},\leq,r)$ be a closed triple satisfying A1-A4. A set $\mathcal{H}\subseteq\mathcal{R}$ is a coideal if it satisfies the following:
  1. (1)
    (Upward-Closure) If $A\in\mathcal{H}$ and $A\leq B$, then $B\in\mathcal{H}$.
  2. (2)
    (A3 mod $\mathcal{H}$) For all $A\in\mathcal{H}$ and $a\in\mathcal{AR}\mathord{\upharpoonright}A$, we have that:
    1. (a)
      $a,B]\cap\mathcal{H}\neq\emptyset$ for all $B\in[0pt_{A}(a),A]\cap\mathcal{H}$.
    2. (b)
      If $B\in\mathcal{H}\mathord{\upharpoonright}A$ and $a,B]\neq\emptyset$, then there exists some $A^{\prime}\in[0pt_{A}(a),A]\cap\mathcal{H}$ such that $\emptyset\neq[a,A^{\prime}]\subseteq[a,B]$.
  3. (3)
    (A4 mod $\mathcal{H}$) Fix $A\in\mathcal{H}$ and $a\in\mathcal{AR}\mathord{\upharpoonright}A$. For any $\mathcal{O}\subseteq\mathcal{AR}_{\operatorname{lh}(a)+1}$, there exists $B\in[0pt_{A}(a),A]\cap\mathcal{H}$ such that $r_{\operatorname{lh}(a)+1}[a,B]\subseteq\mathcal{O}$ or $r_{\operatorname{lh}(a)+1}[a,B]\subseteq\mathcal{O}^{c}$.
A notion of semiselectivity was also given as Definition 3.4-3.7 in [DMN15]. We propose a slight variant of this definition.
Definition 6.3.
Let $(\mathcal{R},\leq,r)$ be an A2-space, and let $\mathcal{H}\subseteq\mathcal{R}$ be a semicoideal. Let $A\in\mathcal{H}$ and $a\in\mathcal{AR}\mathord{\upharpoonright}A$.
  1. (1)
    A family of subsets $\vec{\mathcal{D}}=\{\mathcal{D}_{b}\}_{b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]}$ is dense open below $[a,A]$ in $\mathcal{H}$ if for all $b\in\mathcal{AR}\mathord{\upharpoonright}[b,A]$, $\mathcal{D}_{b}$ is a $\leq$-downward closed subset of $\mathcal{H}\mathord{\upharpoonright}[b,A]$, and for all $B\in\mathcal{H}\mathord{\upharpoonright}[b,A]$, there exists some $C\in\mathcal{H}\mathord{\upharpoonright}[b,B]$ such that $C\in\mathcal{D}_{b}$.
  2. (2)
    Let $\vec{\mathcal{D}}=\{\mathcal{D}_{b}\}_{b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]}$ be dense open below $[a,A]$ in $\mathcal{H}$. We say that $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ diagonalises $\vec{\mathcal{D}}$ if for all $b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]$, there exists some $A_{b}\in\mathcal{D}_{b}$ such that $b,B]\subseteq[b,A_{b}]$.
  3. (3)
    We say that $\mathcal{H}$ is semiselective if for all $A\in\mathcal{H}$, $a\in\mathcal{AR}\mathord{\upharpoonright}A$, every dense open family below $[a,A]$ in $\mathcal{H}$ has a diagonalisation in $\mathcal{H}$.
Lemma 6.4.
Let $(\mathcal{R},\leq,r)$ be a closed triple satisfying A1-A4. Then $\mathcal{R}$ is a semiselective coideal.
This is stated as Lemma 3.11 of [Y24]. We replicate the proof here.
Proof.
Note that $\mathcal{R}$ is a coideal by definition. Fix some $A\in\mathcal{R}$ and $a\in\mathcal{AR}\mathord{\upharpoonright}A$. Suppose that $\vec{\mathcal{D}}=\{\mathcal{D}_{b}\}_{b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]}$ is dense open below $[a,A]$. We shall define a fusion sequence $(A_{n})_{n<\omega}$ in $[a,A]$, with $a_{n+1}:=r_{\operatorname{lh}(a)+n+1}(A_{n})$, such that $A_{n+1}\in[a_{n+1},A_{n}]$: Let $A_{0}:=A$, and suppose that $A_{n}$ has been defined. Let $\{b_{i}:i<N\}$ enumerate the set of all $b\in\mathcal{AR}\mathord{\upharpoonright}A_{n}$ such that $a\sqsubseteq b$ and $b\leq_{\mathrm{fin}}a_{n+1}$. Let $A_{n+1}^{0}:=A_{n}$. If $A_{n+1}^{i}\in[a_{n+1},A_{n}]$ has been defined, let $B_{n+1}\in\mathcal{D}_{b_{i}}$ be such that $B_{n+1}\in[b_{i},A_{n+1}^{i}]$, which exists as $\mathcal{D}_{b_{i}}$ is dense open in $b_{i},A]$. By A3, we then let $A_{n+1}^{i+1}\in[a_{n+1},A_{n}^{i}]$ such that $b_{i},A_{n+1}^{i+1}]\subseteq[b_{i},B_{n+1}]$. We complete the induction by letting $A_{n+1}:=A_{n+1}^{N}$. Let $B$ be the limit of the fusion sequence $(A_{n})_{n<\omega}$, and we have that $B$ diagonalises $\vec{\mathcal{D}}$. ∎
We will see that the semiselectivity property plays an important role in both combinatorial forcing and the abstract Mathias forcing.
6.2. Combinatorial forcing
Let $(\mathcal{R},\leq,r)$ be a closed triple satisfying A1-A4, and let $\mathcal{H}\subseteq\mathcal{R}$ be a coideal. We also fix a subset $\mathcal{X}\subseteq\mathcal{R}$. Unless stated otherwise, we assume that $\mathcal{H}$ is semiselective.
Definition 6.5.
Let $A\in\mathcal{H}$ and let $a\in\mathcal{AR}$. We say that:
  1. (1)
    AA accepts $a$ if $a,A]\subseteq\mathcal{X}$.
  2. (2)
    AA rejects $a$ if $a\in\mathcal{AR}\mathord{\upharpoonright}A$, and for all $\in\mathcal{H}\mathord{\upharpoonright}[a,A], BB$ does not accept $a$.
  3. (3)
    AA decides $a$ if $A$ accepts or rejects $a$.
This localised variant of combinatorial forcing for topological Ramsey spaces was introduced in [DMN15], along with various basic properties of this combinatorial forcing. The details of the proofs of these properties with our definition of semiselectivity (Definition 6.3) remains mostly unchanged, but we shall provide an overview for the sake of completion.
Lemma 6.6.
Let $A\in\mathcal{H}$ and let $a\in\mathcal{AR}$.
  1. (1)
    If $a\notin\mathcal{AR}\mathord{\upharpoonright}A$, $A$ accepts $a$.
  2. (2)
    If $A$ accepts $a$ and $B\in\mathcal{H}\mathord{\upharpoonright}A$, then $B$ accepts $a$.
  3. (3)
    If $A$ rejects $a$, $B\in\mathcal{H}\mathord{\upharpoonright}A$ and $a\in\mathcal{AR}\mathord{\upharpoonright}B$, then $B$ rejects $a$.
  4. (4)
    If $A$ decides $a$ and $B\in\mathcal{H}\mathord{\upharpoonright}A$, then $B$ decides $a$.
  5. (5)
    If $a\in\mathcal{AR}\mathord{\upharpoonright}A$, then there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[0pt_{A}(a),A]$ which decides $B$.
  6. (6)
    If $a\in\mathcal{AR}\mathord{\upharpoonright}A$ and $A$ decides $a$, then for all $\in\mathcal{H}\mathord{\upharpoonright}[0pt_{A}(a),A], BB$ decides $a$ in the same way $A$ does.
  7. (7)
    AA accepts $a$ iff $A$ accepts every $b\in r_{\operatorname{lh}(a)+1}[a,A]$.
Proof.
The proofs are straightforward and mostly the same as the non-localised variant of combinatorial forcing – we refer readers to Lemma 4.31 of [T10]. ∎
Lemma 6.7.
For all $A\in\mathcal{H}$ and $a\in\mathcal{AR}\mathord{\upharpoonright}A$, then there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ which decides every $b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]$.
Proof.
For each $b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]$, let:
$\mathcal{D}_{b}:=\{B\in\mathcal{H}\mathord{\upharpoonright}[b,A]:B\text{ decides }b\}.$
By Lemma 6.6(5), $\mathcal{D}_{b}$ is dense open in $\mathcal{H}\mathord{\upharpoonright}[b,A]$. Since $\mathcal{H}$ is semiselective, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ which diagonalises $\{\mathcal{D}_{b}\}_{b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]}$. If $b\in\mathcal{AR}\mathord{\upharpoonright}B$, then $b,B]\subseteq[b,A_{b}]$ for some $A_{b}\in\mathcal{D}_{b}$. Since $A_{b}$ decides $b$, $B\in\mathcal{H}\mathord{\upharpoonright}A_{b}$ and $b\in\mathcal{AR}\mathord{\upharpoonright}B$, by Lemma 6.6(2) and (3), $B$ decides $b$ as well. Note that if $b\notin\mathcal{AR}\mathord{\upharpoonright}B$, then $b,B]=\emptyset$, so $B$ trivially decides $b$. ∎
Lemma 6.8.
If $A$ rejects $a$, then there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ which rejects every $b\in r_{\operatorname{lh}(a)+1}[a,A]$.
Proof.
By Lemma 6.7, we may let $A^{\prime}\in\mathcal{H}[a,A]$ decide every $b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]$. Let:
$\mathcal{O}:=\{b\in r_{\operatorname{lh}(a)+1}[a,A]:A^{\prime}\text{ rejects }b\}.$
By A4 mod $\mathcal{H}$, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $r_{\operatorname{lh}(a)+1}[a,B]\subseteq\mathcal{O}$ or $r_{\operatorname{lh}(a)+1}[a,B]\subseteq\mathcal{O}^{c}$. It remains to show that the latter case is impossible. Since $A^{\prime}$ decides every $b\in\mathcal{AR}\mathord{\upharpoonright}A$, so does $B$. Therefore, if $r_{\operatorname{lh}(a)+1}[a,B]$ then $B$ accepts every $b\in r_{\operatorname{lh}(a)+1}[a,B]$. By Lemma 6.6(7), this implies that $B$ accepts $a$, contradicting that $A$ rejects $a$. ∎
Lemma 6.9.
If $A$ rejects $a$, then there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ which rejects every $b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]$.
Proof.
Again, by Lemma 6.7 we may let $A^{\prime}\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ decide every $b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]$. For each $b\in\mathcal{AR}\mathord{\upharpoonright}[b,A]$, we let:
$\mathcal{D}_{b}:=\{B\in\mathcal{H}\mathord{\upharpoonright}[b,A]:\text{$B$ accepts $b$ or $B$ rejects every $b^{\prime}\in r_{\operatorname{lh}(b)+1}[b,A]$}\}.$
By Lemma 6.8, $\mathcal{D}_{b}$ is dense open in $\mathcal{H}\mathord{\upharpoonright}[a,A]$, so by the semiselectivity property of $\mathcal{H}$ there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ which diagonalises $\{\mathcal{D}_{b}\}_{b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]}$. We shall induct on $\operatorname{lh}(b)$ and show that $B$ rejects every $b\in\mathcal{AR}\mathord{\upharpoonright}[a,B]$ (if $b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]$ but $b\notin\mathcal{AR}\mathord{\upharpoonright}[a,B]$, then $B$ trivially rejects $b$).
For the base case $\operatorname{lh}(b)=\operatorname{lh}(a)$ (i.e. $b=a$), it follows from that $A$ rejects $a$, and that $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ (cf. Lemma 6.6(3)). If $B$ rejects some $b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]$ where $\operatorname{lh}(b)=n\geq\operatorname{lh}(a)$, then since $b,B]\subseteq[b,A_{b}]$ for some $A_{b}\in\mathcal{D}_{b}$, we have that $B$ rejects every $b^{\prime}\in r_{\operatorname{lh}(a)+1}[b,B]$, completing the induction step. Therefore, $B$ rejects every $b\in\mathcal{AR}\mathord{\upharpoonright}[a,B]$. ∎
The abstract combinatorial forcing allows us to prove the abstract local Ellentuck theorem – see Theorem 3.12 of [DMN15]. We do not require the full strength of this result in this paper. Instead, we give an outline of the proof if $\mathcal{H}$ is semiselective, then every metrically open set $\mathcal{X}\subseteq\mathcal{R}$ is $\mathcal{H}$-Ramsey.
Definition 6.10.
A set $\mathcal{X}\subseteq\mathcal{R}$ is $\mathcal{H}$-Ramsey if for all $A\in\mathcal{R}$ and $a\in\mathcal{AR}\mathord{\upharpoonright}A$, there exists some $B\in[a,A]$ such that $a,B]\subseteq\mathcal{X}$ or $a,B]\subseteq\mathcal{X}^{c}$.
Proposition 6.11.
Let $\mathcal{H}\subseteq\mathcal{R}$ be a semiselective coideal. If $\mathcal{X}\subseteq\mathcal{R}$ is metrically open, then $\mathcal{X}$ is $\mathcal{H}$-Ramsey.
Proof.
Let $A\in\mathcal{H}$ and let $a\in\mathcal{AR}\mathord{\upharpoonright}A$. We consider combinatorial forcing w.r.t. $\mathcal{X}$. If $A$ does not reject $a$, then $a,B]\subseteq\mathcal{X}$ for some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$, so we’re done. Otherwise, $A$ rejects $a$, so by Lemma 6.9 there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $B$ rejects every $b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]$. We claim that $a,B]\subseteq\mathcal{X}^{c}$. Otherwise, we may let $C\in[a,B]\cap\mathcal{X}$. Since $\mathcal{X}$ is metrically open, there exists some $b\in\mathcal{AR}\mathord{\upharpoonright}[a,B]$ such that $C\in[b]\subseteq\mathcal{X}$, where $b]:=\{D\in\mathcal{R}:b\sqsubseteq D\}$. In particular, we have that $b,B]\subseteq\mathcal{X}$, contradicting that $B$ rejects $b$. ∎
6.3. Abstract Mathias forcing
The abstract Mathias forcing localised to a coideal $\mathcal{H}$ is defined as the forcing poset $(\mathbb{M}_{\mathcal{H}},\leq)$, where:
$\mathbb{M}_{\mathcal{H}}=\{(a,A):A\in\mathcal{R}\text{ and }a\in\mathcal{AR}\mathord{\upharpoonright}A\},$
and that $(b,B)\leq(a,A)$ iff $b,B]\subseteq[a,A]$. We also write $(b,B)\leq_{0}(a,A)$ if $(b,B)\leq(a,A)$ and $b=a$.
Recall that given a closed triple $(\mathcal{R},\leq,r)$ satisfying A1-A4, $\mathcal{R}$ is a closed subset of the product space $\mathcal{AR}^{\mathbb{N}}$, where $\mathcal{AR}$ is equipped with the discrete topology. Thus, there exists some absolute formula $\varphi$ such that:
$\mathcal{R}=\{A\in\mathcal{AR}^{\mathbb{N}}:\varphi(A)\}.$
Therefore, for any model $M$, we may write:
$\mathcal{R}^{M}:=\{A\in\mathcal{AR}^{\mathbb{N}}:\varphi^{M}(A)\}.$
Definition 6.12.
An element $A\in\mathcal{R}$ is said to be $\mathbb{M}_{\mathcal{H}}$-generic (over a ground model $\mathbf{V}$) if:
$G_{A}:=\{(a,B)\in\mathbb{M}_{\mathcal{H}}:A\in[a,B]\}$
is a generic filter of $\mathbb{M}_{\mathcal{H}}$ (over $\mathbf{V}$).
Lemma 6.13.
Let $\mathbb{M}_{\mathcal{H}}$ be the Mathias forcing localised to a coideal $\mathcal{H}$.
  1. (1)
    If $G$ is a generic filter of $\mathbb{M}_{\mathcal{H}}$, then there exists a unique $\mathbb{M}_{\mathcal{H}}$-generic $A_{G}\in\mathcal{R}$ such that $G=\{(a,A)\in\mathbb{M}_{\mathcal{H}}:A_{G}\in[a,A]\}$.
  2. (2)
    If $A\in\mathcal{R}$, then $A_{G_{A}}=A$.
  3. (3)
    If $G$ is a generic filter of $\mathbb{M}_{\mathcal{H}}$, then $G_{A_{G}}=G$.
Proof.
We only prove (1) - (2) follows from the uniqueness of $A_{G}$ in (1), and (3) follow easily from the definitions. Let $\mathcal{D}_{n}:=\{(a,A)\in\mathbb{M}_{\mathcal{H}}:\operatorname{lh}(a)\geq n\}$. We see that $\mathcal{D}_{n}$ is dense open in $\mathbb{M}_{\mathcal{H}}$ for all $n$ - if $(a,A)\notin\mathcal{D}_{n}$ (i.e. $\operatorname{lh}(a)<n$) and $b\in r_{n}[a,A]$, then by A3 mod $\mathcal{H}$ we have that $b,A]\cap\mathcal{H}\neq\emptyset$. Let $B\in[b,A]\cap\mathcal{H}$, and we have that $b,B]\subseteq[a,A]$, so $(b,B)\leq(a,A)$ and $(b,B)\in\mathcal{D}_{n}$. Therefore, $G\cap\mathcal{D}_{n}\neq\emptyset$ for all $n$, and since $G$ is a filter, for all $(a,A),(b,B)\in G$, $a\sqsubseteq b$ or $b\sqsubseteq a$. Thus, there exists a unique $A_{G}\in\mathcal{R}$ such that for all $n$, $(r_{n}(A_{G}),A)\in G$ for some $A$. It’s easy to verify that $G=G_{A_{G}}$. ∎
Definition 6.14.
Let $(\mathcal{R},\leq,r)$ be a closed triple satisfying A1-A4, and let $\mathcal{H}\subseteq\mathcal{R}$ be a coideal.
  1. (1)
    $\mathbb{M}_{\mathcal{H}}$ satisfies the Mathias property if whenever $A\in\mathcal{R}$ is $\mathbb{M}_{\mathcal{H}}$-generic over a ground model $\mathbf{V}$ and $\leq A, BB$ is also $\mathbb{M}_{\mathcal{H}}$-generic over $\mathbf{V}$.
  2. (2)
    $\mathbb{M}_{\mathcal{H}}$ satisfies the Prikry property if whenever $\varphi$ is a sentence in the forcing language, then we have that for all $p\in\mathbb{M}_{\mathcal{H}}$, there exists some $q\leq_{0}p$ such that $q\Vdash\varphi$ or $q\Vdash\neg\varphi$.
Given a subset $\mathcal{D}\subseteq\mathbb{M}_{\mathcal{H}}$, we let:
$\overline{\mathcal{D}}$ $:=\bigcup_{(a,A)\in\mathcal{D}}[a,A].$
Note that $\overline{\mathcal{D}}$ is an Ellentuck-open subset of $\mathcal{R}$ in $\mathbf{V}$. Thus, if $\mathcal{D}$ is dense open in $\mathbb{M}_{\mathcal{H}}$, then for all $A\in\mathcal{H}$ and $a\in\mathcal{AR}\mathord{\upharpoonright}A$, $A$ does not reject $a$ w.r.t. $\overline{\mathcal{D}}$ (otherwise, by Lemma 6.9 there exists some $B\in[a,A]$ which rejects every $b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]$, contradicting that $\mathcal{D}$ is dense in $\mathbb{M}_{\mathcal{H}}$). In other words, for all $A\in\mathcal{H}$ and $a\in\mathcal{AR}\mathord{\upharpoonright}A$, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $a,B]\subseteq\overline{\mathcal{D}}$.
Notation 6.15.
Given $a,b\in\mathcal{AR}$, we write:
$b\preceq_{\mathrm{fin}}a\iff b\leq_{\mathrm{fin}}a\text{ and }\forall a^{\prime}\sqsubseteq a\,(a^{\prime}\neq a\to b\not\leq_{\mathrm{fin}}a^{\prime}).$
Proposition 6.16.
If $\mathcal{H}$ is semiselective, then $\mathbb{M}_{\mathcal{H}}$ has the Mathias property.
Proof.
Let $A_{G}$ be an $\mathbb{M}_{\mathcal{H}}$-generic element of $\mathcal{R}$ (over some ground model $\mathbf{V}$), and let $A_{H}\leq A_{G}$. We need to show that the filter:
$H:=\{(a,A)\in\mathbb{M}_{\mathcal{H}}:A_{H}\in[a,A]\}$
is generic. Let $\mathcal{D}\subseteq\mathbb{M}_{\mathcal{H}}$ be a dense open subset (in the ground model $\mathbf{V}$). For each $a\in\mathcal{AR}$, we let:
$\mathcal{E}_{a}:=\{B\in\mathcal{H}:a\in\mathcal{AR}\mathord{\upharpoonright}B\text{ and }\forall b\preceq_{\mathrm{fin}}a\,[b,B]\subseteq\overline{\mathcal{D}}\}.$
Claim 10.
For all $A\in\mathcal{H}$ such that $\in\mathcal{AR}\mathord{\upharpoonright}A, a\mathcal{E}_{a}$ is dense open in $\mathcal{H}\mathord{\upharpoonright}[a,A]$.
Proof.
It’s clear that $\mathcal{E}_{a}$ is open, so shall show that it is dense in $\mathcal{H}\mathord{\upharpoonright}[a,A]$. By the comment above, let $A^{\prime}\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $a,A^{\prime}]\subseteq\overline{\mathcal{D}}$. Let $\{b_{i}:1\leq i\leq N\}$ be an enumeration of all $b\preceq_{\mathrm{fin}}a$. Let $A_{0}:=A^{\prime}$, and suppose that we have defined $A_{i}\in[a,A^{\prime}]$. We have that $b_{i+1}\in\mathcal{AR}\mathord{\upharpoonright}A_{i}$ (as $b_{i+1}\leq_{\mathrm{fin}}a\sqsubseteq A_{i}$), so by the observation stated prior to the claim there exists some $A_{i+1}^{\prime}\in\mathcal{H}\mathord{\upharpoonright}[b_{i+1},A_{i}]$ which accepts $b_{i+1},A_{i}$ w.r.t. $\overline{\mathcal{D}}$, i.e. $b_{i+1},A_{i+1}^{\prime}]\subseteq\overline{\mathcal{D}}$. By A3 mod $\mathcal{H}$, there exists some $A_{i+1}\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $a_{i+1},A_{i+1}]\subseteq[a_{i+1},A_{i+1}^{\prime}]$. This completes the induction, and $(a,A_{N})\in\mathcal{D}^{\prime}$, as desired. ∎
Fix any $(a,A)\in G$. Since $\mathcal{H}$ is semiselective, for all $(b,B)\leq(a,A)$ in $\mathbb{M}_{\mathcal{H}}$, there exists some $C\in\mathcal{H}\mathord{\upharpoonright}[b,B]$ which diagonalises $\vec{\mathcal{E}}_{b}:=\{\mathcal{E}_{c}\}_{c\in\mathcal{AR}\mathord{\upharpoonright}[b,A]}$ within $A$. That is, the set:
$\mathcal{D}^{\prime}:=\{(b,B)\in\mathbb{M}_{\mathcal{H}}:\text{$B$ diagonalises $\vec{\mathcal{E}}_{b}$ within $A$}\}$
is dense open in $\mathbb{M}_{\mathcal{H}}$. Since $G$ is generic, we may let $(b,B)\in G\cap\mathcal{D}^{\prime}$. Then $A_{G}\in[b,B]$, and for all $c\in\mathcal{AR}\mathord{\upharpoonright}[b,A_{G}]$ and $d\preceq_{\mathrm{fin}}c$, $d,A_{G}]\subseteq[d,B]\subseteq\overline{\mathcal{D}}$. Let $n$ be the least integer such that $0pt_{A_{G}}(r_{n}(A_{H}))=:m\geq\operatorname{lh}(b)$. Then $r_{n}(A_{H})\preceq_{\mathrm{fin}}r_{m}(A_{G})\in\mathcal{AR}\mathord{\upharpoonright}[b,B]$, so $A_{H}\in[r_{n}(A_{H}),A_{G}]\subseteq[r_{n}(A_{H}),B]\subseteq\overline{\mathcal{D}}$. This implies that $(r_{n}(A_{H}),B)\in H\cap\mathcal{D}$, and since $\mathcal{D}$ is arbitrary, $H$ is a generic filter as required. ∎
Proposition 6.17.
If $\mathcal{H}$ is semiselective, then $\mathbb{M}_{\mathcal{H}}$ has the Prikry property.
Proof.
Let $\varphi$ be a sentence in the forcing language. Let:
$\mathcal{D}$ $:=\{q\in\mathbb{M}_{\mathcal{H}}:q\Vdash\varphi\}.$
We consider doing combinatorial forcing w.r.t. the set $\overline{\mathcal{D}}$. Let $(a,A)\in\mathbb{M}_{\mathcal{H}}$, and by Lemma 6.6(2) and (3) there exists some $B\in[a,A]$ which decides $a$. If $B$ accepts $a$, then $a,B]\subseteq\overline{\mathcal{D}}$, so $(a,B)\in\mathcal{D}$ as $\mathcal{D}$ is open. In this case, $(a,B)\leq_{0}(a,A)$ and $(a,B)\Vdash\varphi$. Otherwise, by Lemma 6.9 there exists some $B\in[a,A]$ which rejects every $b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]$. In other words, for all $(b,C)\leq(a,B)$, $(b,C)\notin\mathcal{D}_{0}$, i.e. $(b,C)$ does not force $\varphi$. But this precisely means that $(a,B)\Vdash\neg\varphi$, as needed. ∎
7. Topological Ramsey space of block subspaces
By Hindman’s theorem, the pigeonhole principle hold for $E^{[\infty]}$ if $|\mathbb{F}|=2$. That is, for all $Y\subseteq E$ and $A\in E^{[\infty]}$, there exists some $B\leq A$ such that $\langle B\rangle\subseteq Y$ or $\langle B\rangle\subseteq Y^{c}$. Consequently, if for each $A:=(x_{n})_{n<\omega}\in E^{[\infty]}$ we let:
$r_{n}(A):=(x_{0},\dots,x_{n-1}),$
then $(E^{[\infty]},\leq,r)$ is a closed triple satisfying A1-A4. We shall discuss the intersection of the local Ramsey theory of $(E^{[\infty]},\leq,r)$ as a topological Ramsey space, and the local Ramsey theory of block subspaces developed Smythe, followed by an application of the abstract Mathias forcing to prove Theorem 1.7.
7.1. Semicoideals of block subspaces
Readers may notice that there are two separate definitions for a coideal $\mathcal{H}\subseteq E^{[\infty]}$, i.e. Definition 3.4 and Definition 6.2. It turns out that these two definitions are equivalent. We remark that in this setting, if $A\in E^{[\infty]}$ and $a\in E^{[<\infty]}\mathord{\upharpoonright}A$ (i.e. $a\in E^{[<\infty]}$ and $\langle a\rangle\subseteq\langle A\rangle$), we have that $0pt_{A}(a)=d_{A}(a)$ (see Notation 2.2).
Lemma 7.1.
Let $\mathcal{H}\subseteq E^{[\infty]}$. The following are equivalent:
  1. (1)
    $\mathcal{H}$ is a semicoideal.
  2. (2)
    $\mathcal{H}$ is $\leq$-upward closed and A3 mod $\mathcal{H}$ holds.
Proof.
On one hand, if $A\in\mathcal{H}$ and $a,A]\subseteq[a,B]$ for some $a\in\mathcal{AR}\mathord{\upharpoonright}A$, then by A3 mod $\mathcal{H}$ there exists some $C\in[a,A]\cap\mathcal{H}$. Then $C\in[a,B]$, and in particular $C\leq B$, so $B\in\mathcal{H}$ by $\leq$-upward closure. Conversely, suppose that $\mathcal{H}$ is $\leq^{*}$-upward closed. Let $A\in\mathcal{H}$ and $a\in E^{[<\infty]}\mathord{\upharpoonright}A$.
  1. (1)
    For any $B\in[d_{A}(a),A]\cap\mathcal{H}$, we have that $B\leq^{*}B/a$ (as $B/a\leq B/a$), so $B/a\in\mathcal{H}$. Since $B/a\leq a^{\frown}(B/a)$, we have that $a^{\frown}(B/a)\in\mathcal{H}$ by $\leq$-upward closure. But $a^{\frown}(B/a)\leq^{*}B$, so $a^{\frown}(B/a)\in[a,B]\cap\mathcal{H}$ as required.
  2. (2)
    Suppose that $B\in\mathcal{H}\mathord{\upharpoonright}A$ and $a,B]\neq\emptyset$. Let $N:=d_{A}(a)$, and let $M$ be large enough so that $B/M\leq A/N$. Then $C:=r_{n}(A)^{\frown}(B/m)\in[d_{A}(a),A]\cap\mathcal{H}$ is well-defined, and:
    $D\in[a,C]$ $\implies a\sqsubseteq D\wedge D\leq C$
    $\implies a\sqsubseteq D\wedge D/a\leq C/a$
    $\implies a\sqsubseteq D\wedge D/a\leq B/M$
    $\implies a\sqsubseteq D\wedge D/a\leq A/N$
    $\implies D\in[a,A],$
    as desired.
Lemma 7.2.
Let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal. Then $\mathcal{H}$ is a coideal (as in Definition 3.4) iff A4 mod $\mathcal{H}$ holds.
Proof.
Definition 3.4 is exactly the statement A4 mod $\mathcal{H}$ for $a=\emptyset$, so it suffices to show that the converse holds. Let $A\in E^{[\infty]}$, $a\in E^{[<\infty]}\mathord{\upharpoonright}A$, and let $\mathcal{O}\subseteq E^{[\operatorname{lh}(a)+1]}$. Define:
$Y:=\{x\in E:a<x\wedge a^{\frown}x\in\mathcal{O}\}.$
Since $\mathcal{H}$ is a coideal (as in Definition 3.4), there exists some $B^{\prime}\leq A$ such that $\langle B^{\prime}\rangle\subseteq Y$ or $\langle B^{\prime}\rangle\subseteq Y^{c}$. Let $M$ be large enough so that $B^{\prime}/M\leq A/d_{A}(a)$, and note that $a<B^{\prime}/M$. Let $B:=B^{\prime}/M$, and let $C:=a^{\frown}B\in[a,A]$. Observe that:
$b\in r_{\operatorname{lh}(a)+1}[a,C]\iff b=a^{\frown}x\text{ for some $x\in\langle B\rangle$}.$
Therefore, if $\langle B\rangle\subseteq Y$, then $r_{\operatorname{lh}(a)+1}[a,C]\subseteq\mathcal{O}$, and if $\langle B\rangle\subseteq Y^{c}$, then $r_{\operatorname{lh}(a)+1}[a,C]\subseteq\mathcal{O}^{c}$. Thus, A4 mod $\mathcal{H}$ holds. ∎
We now address the overlap in the selectivity properties. There are two natural questions:
  1. (1)
    Is Definition 6.3 equivalent to the natural extension of semiselectivity of coideal of $\omega]^{\omega}$ to semicoideal of block subspaces (i.e. Lemma 7.10, [T10])?
  2. (2)
    Is a selective (semi)coideal semiselective?
We provide a positive answer to both of these problems.
Lemma 7.3.
Let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal. The following are equivalent:
  1. (1)
    $\mathcal{H}$ is semiselective (as in Definition 6.3).
  2. (2)
    If $\{\mathcal{D}_{n}\}_{n<\omega}$ is a collection of dense open subsets of $\mathcal{H}$ (under $\leq$), then for all $A\in\mathcal{H}$ there exists some $B\in\mathcal{H}\mathord{\upharpoonright}A$ such that $B/n\in\mathcal{D}_{d_{A}(r_{n}(B))}$ for all $n$.
In this statement (2) of the lemma, we also say that $B$ diagonalises $\{\mathcal{D}_{n}\}_{n<\omega}$ within $A$.
Proof.
[…]
$\mathcal{D}_{b}:=\{b^{\frown}(B/b):B\in\mathcal{D}_{d_{A}(b)}\text{ and }B\leq A\}.$
It’s clear that $\mathcal{D}_{b}\subseteq\mathcal{H}\mathord{\upharpoonright}[b,A]$ for all $b\in\mathcal{AR}\mathord{\upharpoonright}A$, and that $\mathcal{D}_{b}$ is $\leq$-downward closed in $\mathcal{H}\mathord{\upharpoonright}[b,A]$. We see that $\mathcal{D}_{b}$ is dense in $\mathcal{H}\mathord{\upharpoonright}[b,A]$ – if $B\in\mathcal{H}\mathord{\upharpoonright}[b,A]$, then since $\mathcal{D}_{d_{A}(b)}$ is dense open in $\mathcal{H}\mathord{\upharpoonright}A$, there exists some $C\in\mathcal{D}_{d_{A}(b)}$ such that $C\leq B$. Then $b^{\frown}(C/b)\in\mathcal{D}_{b}$ and $b^{\frown}(C/b)\leq B$. Since $\mathcal{H}$ is semiselective, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}A$ such that for all $b\in\mathcal{AR}\mathord{\upharpoonright}B$, $b,B]\subseteq[b,A_{b}]$ for some $A_{b}\in\mathcal{D}_{b}$. In particular, $B\in[r_{n}(B),B]\subseteq[r_{n}(B),A_{r_{n}(B)}]$ for some $A_{r_{n}(B)}\in\mathcal{D}_{r_{n}(B)}$, so $B/n\leq A_{r_{n}(B)}/r_{n}(B)\in\mathcal{D}_{d_{A}(r_{n}(B))}$.
[…]
$\mathcal{D}_{b}^{\prime}$ $:=\{B/b:B\in\mathcal{D}_{b}\},$
$\mathcal{D}_{n}$ $:=\bigcap_{\begin{subarray}{c}b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]\\ d_{A}(b)=d_{A}(a)+n\end{subarray}}\mathcal{D}_{b}^{\prime}.$
We claim that $\mathcal{D}_{b}^{\prime}$ is dense open in $\mathcal{H}\mathord{\upharpoonright}A$ for all $b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]$. This implies that $\mathcal{D}_{n}$ is dense open in $\mathcal{H}\mathord{\upharpoonright}A$, as it is a finite intersection of dense open subsets of $\mathcal{H}\mathord{\upharpoonright}A$. Let $B\in\mathcal{H}\mathord{\upharpoonright}A$, and let $b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]$. Then $b^{\frown}(B/b)\in\mathcal{H}\mathord{\upharpoonright}[b,A]$, and since $\mathcal{D}_{b}$ is dense open in $\mathcal{H}\mathord{\upharpoonright}[b,A]$, there exists some $C\in\mathcal{D}_{b}$ such that $C\leq b^{\frown}(B/b)$. This implies that $C/b\leq B/b$, and $C/b\in\mathcal{D}_{b}^{\prime}$, as desired.
By our assumption, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}A$ such that $B/n\in\mathcal{D}_{d_{A}(r_{n}(B))}$ for all $n$. Observe that $a<B$, as $B\in\mathcal{D}_{0}\subseteq\mathcal{D}_{a}$, so $B\leq C/a$ for some $C\in\mathcal{D}_{a}$. Therefore, $a^{\frown}B\in\mathcal{R}$ is well-defined. We claim that $a^{\frown}B$ diagonalises $\vec{\mathcal{D}}$. Let $b\in\mathcal{AR}\mathord{\upharpoonright}[a,a^{\frown}B]$, and suppose that $d_{A}(b)=d_{A}(a)+n$. Let $b^{\prime}$ be such that $b=a^{\frown}b^{\prime}$, and note that $d_{A}(b^{\prime})=d_{A}(b)$. Since $\langle b^{\prime}\rangle\subseteq\langle r_{d_{B}(b^{\prime})}(B)\rangle$, we have that $d_{A}(b^{\prime})\leq d_{A}(r_{d_{B}(b^{\prime})}(B))$. Therefore:
$(a^{\frown}B)/b$ $=B/b^{\prime}=B/d_{B}(b^{\prime})\in\mathcal{D}_{d_{A}(r_{d_{B}(b^{\prime})}(B))}\subseteq\mathcal{D}_{d_{A}(b^{\prime})}=\mathcal{D}_{d_{A}(b)}\subseteq\mathcal{D}_{b}^{\prime},$
so $A_{b}:=b^{\frown}((a^{\frown}B)/b)\in\mathcal{D}_{b}$. Observe that $b,B]=[b,A_{b}]$, and since $b$ is arbitrary, $B$ diagonalises $\vec{\mathcal{D}}$. ∎
Lemma 7.4.
Let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal. If $\mathcal{H}$ is selective, then $\mathcal{H}$ is semiselective.
Proof.
Let $\{\mathcal{D}_{n}\}_{n<\omega}$ be a family of dense open subsets of $\mathcal{H}$, and let $A\in\mathcal{H}$. Define a $\leq$-decreasing sequence in $\mathcal{H}$ as follows: Let $A_{0}\in\mathcal{D}_{0}$ be such that $A_{0}\leq A$. If $A_{n}\in\mathcal{D}_{n}$ has been defined, then let $A_{n+1}\in\mathcal{D}_{n+1}$ be such that $A_{n+1}\leq A_{n}$. By the selectivity of $\mathcal{H}$, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}A$ such that $B/n\leq A_{d_{A}(r_{n}(B))}\in\mathcal{D}_{d_{A}(r_{n}(B))}$ for all $n$. Since $\mathcal{D}_{n}$’s are all $\leq$-downward closed, $B/n\in\mathcal{D}_{d_{A}(r_{n}(B))}$ for all $n$. ∎
An abstract notion of selectivity was also proposed in [DMN15], but the definition appears to be problematic. For instance, Proposition 5.4 of [DMN15] claims that every selective coideal is semiselective, but there appears to be a gap in the proof – it is unclear how the element $A^{2,1}\in\mathcal{D}_{a_{1}^{2}}\mathord{\upharpoonright}A^{1,n_{1}}$ may be obtained. It seems unlikely that this version of selectivity coincides with our definition of selectivity.
Definition 7.5.
Let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal.
  1. (1)
    $\mathcal{H}$ is a filter if $(\mathcal{H},\leq)$ is a filter (as a partial order), i.e. for all $A,B\in\mathcal{H}$ there exists some $C\in\mathcal{H}$ such that $C\leq A$ and $C\leq B$.
  2. (2)
    $\mathcal{H}$ is an ultrafilter if it is a coideal and a filter.
We shall dedicate the remaining portion of this subsection to showing that if $\mathcal{U}\subseteq^{[\infty]}$ is a selective ultrafilter, then a block sequence $A_{G}\in(E^{[\infty]})^{\mathbf{V}[G]}$ is $\mathbb{M}_{\mathcal{U}}$-generic iff $A_{G}\leq^{*}A$ for all $A\in\mathcal{U}$. The proof is very similar to the case for the Ellentuck space – a proof may be found in, for instance, [H22].
The following proposition, in the setting of topological Ramsey spaces, was stated as Theorem 6.6 of [DMN15], extending the original theorem in the setting $\omega]^{\omega}$ in [F98]. We shall reprove this result for block subspaces.
Proposition 7.6.
Assume that $|\mathbb{F}|=2$. Let $\mathcal{H}\subseteq E^{[\infty]}$ be a coideal. The following are equivalent:
  1. (1)
    $\mathcal{H}$ is semiselective.
  2. (2)
    $(\mathcal{H},\leq^{*})$ is $\sigma$-distributive, and every $(\mathcal{H},\leq^{*})$-generic filter $\mathcal{U}$ is a selective ultrafilter.
Proof.
[…]
Let $\mathcal{U}$ be a $(\mathcal{H},\leq^{*})$-generic filter. It’s clear that $\mathcal{U}$ is a filter (as in Definition 7.5). We need to show that $\mathcal{U}$ is a coideal (and hence an ultrafilter). Let $Y\subseteq E$, and let $A\in\mathcal{U}\subseteq\mathcal{H}$. Since $Y$ is countable and $(\mathcal{H},\leq^{*})$ adds no reals, $Y$ is in the ground model. By Hindman’s theorem, the set:
$\mathcal{D}:=\{B\in\mathcal{H}\mathord{\upharpoonright}A:\langle B\rangle\subseteq Y\text{ or }\langle B\rangle\subseteq Y^{c}\}$
is dense open below $A$ in $(\mathcal{H},\leq^{*})$. Therefore, the genericity of $\mathcal{U}$ implies that $\mathcal{U}\cap\mathcal{D}\neq\emptyset$, i.e. there exists some $B\leq A$ in $\mathcal{U}$ such that $\langle B\rangle\subseteq Y$ or $\langle B\rangle\subseteq Y^{c}$. Hence, $\mathcal{U}$ is a coideal.
We now show that $\mathcal{U}$ is selective. Let $(A_{n})_{n<\omega}$ be a $\leq$-decreasing sequence in $\mathcal{U}$, and let $A:=A_{0}$. Since $(\mathcal{H},\leq^{*})$ adds no reals, the sequence $(A_{n})_{n<\omega}$ is in the ground model. For each $n$, let:
$\mathcal{D}_{n}:=\{B\in\mathcal{H}\mathord{\upharpoonright}A:B\leq A_{n}\text{ or }(\nexists C\in\mathcal{H}\mathord{\upharpoonright}A\text{ s.t. }C\leq B\text{ and }C\leq A_{n})\}.$
It’s clear that $\mathcal{D}_{n}$ is dense open below $A$ in $\mathcal{H}$ for all $n$. Since $\mathcal{H}$ is semiselective, we have that the set:
$\mathcal{E}:=\{B\in\mathcal{H}\mathord{\upharpoonright}A:\text{$B$ diagonalises $\{\mathcal{D}_{n}\}_{n<\omega}$ within $A$}\}$
is dense open below $A$ in $\mathcal{H}$. By the genericity of $\mathcal{U}$, let $B\in\mathcal{U}\cap\mathcal{E}$. Then $B/n\in\mathcal{D}_{d_{A}(r_{n}(B))}$ for all $n$, and since $\mathcal{U}$ is a filter, $B$ is compatible with $A_{n}$ for all $n$ in $\mathcal{U}$, so $B/n\leq A_{d_{A}(r_{n}(B))}$ for all $n$. Thus, $B$ diagonalises $(A_{n})_{n<\omega}$ within $A$.
[…]
Definition 7.7.
Let $(\mathcal{R},\leq,r)$ be a closed triple satisfying A1-A4, and let $\mathcal{H}\subseteq\mathcal{R}$. Let $\mathcal{D}\subseteq\mathbb{M}_{\mathcal{H}}$. We say that $(a,A)\in\mathbb{M}_{\mathcal{H}}$ captures $\mathcal{D}$ if for all $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$, there exists some initial segment $b\in\mathcal{AR}\mathord{\upharpoonright}[a,B]$ of $B$ such that $(b,B)\in\mathcal{D}$.
We remark that if $\mathcal{AR}$ is countable (e.g. $E^{[<\infty]}$ is countable), then the statement “$(a,A)$ captures $\mathcal{D}$” is $\Pi_{1}$ with respect to the Polish topology of $\mathcal{R}$. Since $\mathcal{R}$ is a subspace of $\mathcal{AR}^{\mathbb{N}}$ and $\mathcal{AR}$ is equipped with the discrete topology, we may apply Mostowki’s absoluteness th55eorem to conclude that “$(a,A)$ captures $\mathcal{D}$” is absolute.
Lemma 7.8.
Assume that $|\mathbb{F}|=2$. Let $\mathcal{U}\subseteq E^{[\infty]}$ be a selective ultrafilter. Let $\mathcal{D}\subseteq\mathbb{M}_{\mathcal{U}}$ be dense open. For all $a\in E^{[<\infty]}$ and $A\in E^{[\infty]}$ such that $a\in E^{[<\infty]}\mathord{\upharpoonright}A$, there exists some $B\in\mathcal{U}\mathord{\upharpoonright}[a,A]$ such that $(a,A)$ captures $\mathcal{D}$.
Proof.
For each $b\in E^{[<\infty]}\mathord{\upharpoonright}[a,A]$, let $A_{b}\in\mathcal{U}\mathord{\upharpoonright}[b,A]$ be any block sequence such that $(b,A_{b})\in\mathcal{D}$ if it exists – otherwise, we let $A_{b}:=A$. Since $\mathcal{U}$ is an ultrafilter, we may inductively define a $\leq$-decreasing sequence in $\mathcal{U}\mathord{\upharpoonright}[a,A]$ such that $A_{n}\leq A_{b}/a$ for all $b\in E^{[<\infty]}\mathord{\upharpoonright}[a,A]$ in which $d_{A}(b)\leq n$. Since $\mathcal{U}$ is selective, there exists some $B^{\prime}\in\mathcal{H}\mathord{\upharpoonright}(A/a)$ which diagonalises $(A_{n})_{n<\omega}$ within $A$. Let $B:=a^{\frown}B^{\prime}\in\mathcal{U}\mathord{\upharpoonright}[a,A]$, and consider the set:
$\mathcal{X}:=\{C\in E^{[\infty]}:C\in[a,B]\to(\exists b\in E^{[<\infty]}\mathord{\upharpoonright}[a,C]\text{ s.t. }(b,B)\in\mathcal{D})\}.$
Since $\mathcal{X}$ is a metrically open subset of $E^{[<\infty]}$, by Proposition 6.11 there exists some $C\in\mathcal{U}\mathord{\upharpoonright}[a,B]$ such that $a,C]\subseteq\mathcal{X}$ or $a,C]\subseteq\mathcal{X}^{c}$. Observe that if $a,C]\subseteq\mathcal{X}$, then $(a,C)$ captures $\mathcal{D}$, so we’re done. It remains to show that $a,C]\subseteq\mathcal{X}^{c}$ is not possible. Since $\mathcal{D}$ is dense in $\mathbb{M}_{\mathcal{H}}$, there exists some condition $(b,D)\leq(a,C)$ in $\mathcal{D}$. This means that $(b,A_{b})\in\mathcal{D}$, and since $B/b\leq A_{d_{A}(b)}\leq A_{b}$, we have that $(b,B)\in\mathcal{D}$. In other words, there exists some $b\in E^{[<\infty]}\mathord{\upharpoonright}C$ such that $(b,B)\in\mathcal{D}$, so $C\notin\mathcal{X}$. ∎
Proposition 7.9.
Assume that $|\mathbb{F}|=2$. Let $\mathcal{U}\subseteq E^{[\infty]}$ be a selective ultrafilter. Then a block sequence $A_{G}$ is $\mathbb{M}_{\mathcal{U}}$-generic iff $A_{G}\leq^{*}A$ for all $A\in\mathcal{U}$.
Proof.
Given a block sequence $A_{G}$, we let:
$G:=\{(a,A)\in\mathbb{M}_{\mathcal{U}}:A_{G}\in[a,A]\}.$
$\implies$: Suppose that $A_{G}$ is $\mathbb{M}_{\mathcal{U}}$-generic, i.e. $G$ is a generic filter. For any $A\in\mathcal{U}$ and $a\in E^{[<\infty]}\mathord{\upharpoonright}A$, the set:
$\mathcal{D}_{A}:=\{(b,B)\in\mathbb{M}_{\mathcal{U}}:B\leq^{*}A\text{ and }b\in E^{[<\infty]}\mathord{\upharpoonright}B\}$
is dense open in $\mathbb{M}_{\mathcal{U}}$ – indeed, for any $(b,B)\in\mathbb{M}_{\mathcal{U}}$, let $C\in\mathcal{U}$ be such that $C\leq B$ and $C\leq A$. Then $(a,a^{\frown}(C/a))\in\mathcal{D}_{A}$, and $C\leq^{*}B$ and $C\leq^{*}B$. Since $G$ is generic, $G\cap\mathcal{D}_{A}\neq\emptyset$, so we may let $(b,B)\in\mathcal{D}_{A}$ such that $A_{G}\in[b,B]$. This implies that $A_{G}\leq^{*}B\leq^{*}A$, as desired.
[…]
We have that $A_{G}\leq^{*}B$, so let $a\in E^{[<\infty]}\mathord{\upharpoonright}A_{G}\cap E^{[<\infty]}\cap B$ be such that $A_{G}/a\leq B/a\leq A_{a}$. Therefore, $(a,B)$ captures $\mathcal{D}$ in the ground model, so $(a,B)$ captures $\mathcal{D}$ in the generic extension (see the remark after Definition 7.7). Since $A_{G}\in[a,B]$, we have that $(r_{n}(A_{G}),B)\in\mathcal{D}$ for some $n$, as desired. ∎
7.2. Proof of Theorem 1.7
Recall that for an uncountable cardinal $\kappa$, the forcing poset $\operatorname{Coll}(\omega,<\!\kappa)$ is the Lévy collapse, which collapses $\kappa$ to $\omega_{1}$. We shall follow the proof layout presented in [NN18] (for the proof of Theorem 16). The following lemma is due to Solovay, and a proof may be found in [solovay]. For the rest of this section, we fix a ground model $\mathbf{V}$ of $\mathsf{ZFC}$, and all vector spaces are over $\mathbb{F}_{2}$.
Lemma 7.10.
Let $\kappa$ be an inaccessible cardinal, and let $\mathbb{P}\in\mathbf{V}$ be a forcing poset of size $\kappa$. Let $G$ be a $\operatorname{Coll}(\omega,<\!\kappa)$-generic filter. If $H\in\mathbf{V}[G]$ is $\mathbb{P}$-generic over $\mathbf{V}$, then there exists a filter $G^{*}$ which is $\operatorname{Coll}(\omega,<\!\kappa)$-generic over $\mathbf{V}[H]$, and $\mathbf{V}[H][G^{*}]=\mathbf{V}[G]$.
Proof of Theorem 1.7.
Let $\mathcal{H}\in\mathbf{V}[G]$ be a selective coideal of $E^{[\infty]}$, and let $\mathcal{X}\in\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$ be a subset of $E^{[\infty]}$. Since every set in $\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$ is ordinal-definable from some real in $\mathbf{V}[G]$, there exists a ternary formula $\phi$, some real $r\in\mathbf{V}[G]$ and a sequence of ordinals $\alpha$ such that for all $B\in(E^{[\infty]})^{\mathbf{V}[G]}$:
$B\in\mathcal{X}\iff\mathbf{V}[G]\models\phi[B,r,\alpha].$
We need to show that for all $A\in\mathcal{H}$ and $a\in E^{[<\infty]}\mathord{\upharpoonright}A$, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $a,B]\subseteq\mathcal{X}$ or $a,B]\subseteq\mathcal{X}$. We fix some $A\in\mathcal{H}$.
Since $\kappa$ is Mahlo, we may fix some inaccessible $\lambda<\kappa$ such that $\mathcal{H}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]\in[G\mathord{\upharpoonright}\lambda]$, and $\mathbf{V}[G\mathord{\upharpoonright}\lambda]\models\mathcal{H}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]$ is a selective coideal, and denote $\overline{\mathcal{H}}:=\mathcal{H}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]$. By the $\kappa$-chain condition of $\operatorname{Coll}(\omega,<\!\kappa)$, we may also assume that $A\in\overline{\mathcal{H}}$ and $r\in\mathbf{V}[G\mathord{\upharpoonright}\lambda]$. Let $\phi^{*}$ be the formula such that $\phi^{*}(x,y,w)$ holds iff $\Vdash_{\operatorname{Coll}(\omega,<\kappa)}\phi(\dot{x},\check{y},\check{w})$. Let:
$\overline{\mathcal{X}}:=\{B\in E^{[\infty]}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]:V[G\mathord{\upharpoonright}\lambda]\models\phi[B,r,\alpha]\}.$
Note that $\overline{\mathcal{X}}\in\mathbf{V}[G\mathord{\upharpoonright}\lambda]$. Furthermore, since the Lévy collapse is homogeneous, we have that for all $B\in E^{[\infty]}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]$:
$B\in\overline{\mathcal{X}}$ $\iff\mathbf{V}[G]\models\phi[B,r,\alpha]$
$\iff\Vdash_{\operatorname{Coll}(\omega,<\kappa)}\phi[B,r,\alpha]$
$\iff B\in\overline{\mathcal{X}}.$
Thus, $\overline{\mathcal{X}}=\mathcal{X}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]$.
Let $\mathcal{U}$ be a $(\overline{\mathcal{H}},\leq^{*})$-generic filter such that $A\in\mathcal{U}$. Note that by Proposition 7.6, $\mathcal{U}$ is a selective ultrafilter. Since $(\overline{\mathcal{H}},\leq^{*})$ adds no reals, we have that $E^{[\infty]}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]=E^{[\infty]}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$. This implies that $\overline{\mathcal{X}}=\mathcal{X}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$. Furthermore, we have that:
$\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}][G\mathord{\upharpoonright}[\lambda,\kappa)]=\mathbf{V}[G\mathord{\upharpoonright}\lambda][G\mathord{\upharpoonright}[\lambda,\kappa)][\mathcal{U}]=\mathbf{V}[G][\mathcal{U}].$
Claim 11.
Let $\mathbb{P}\in\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$ be some forcing notion which adds an element of $E^{[\infty]}$, and let $B\in E^{[\infty]}\cap\mathbf{V}[G][\mathcal{U}]$ be $\mathbb{P}$-generic over $\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$. Then:
$B\in\mathcal{X}\iff V[G\mathord{\upharpoonright}\lambda][\mathcal{U}][B]\models\phi^{*}[B,r,\alpha].$
Proof.
We work in $\mathbf{V}[G][\mathcal{U}]$. Since $\kappa$ is inaccessible, $\operatorname{Coll}(\omega,<\!\lambda)*\dot{\mathbb{P}}$ is a poset in $\mathbf{V}$ of size $\kappa$, so by Lemma 7.10 applied to the model $\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$ there exists a filter $G^{*}$, $\operatorname{Coll}(\omega,<\!\kappa)$-generic over $\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}][B]$ such that:
$\mathbf{V}[G][\mathcal{U}]=\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}][G\mathord{\upharpoonright}[\lambda,\kappa)]=\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}][B][G^{*}].$
Therefore:
$B\in\mathcal{X}$ $\iff\mathbf{V}[G][\mathcal{U}]\models\phi[B,r,\alpha]$
$\iff\;\Vdash\phi(\dot{B},\check{r},\check{\alpha})\text{ in $\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}][B]$}$
$\iff\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}][B]\models\phi^{*}[B,r,\alpha],$
where the second equivalence follows from the homogeneity of the Lévy collapse. ∎
We consider the forcing notion $\mathbb{M}_{\mathcal{U}}$ in $\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$, and note that $(a,A)\in\mathbb{M}_{\mathcal{U}}$. By the Prikry property of $\mathbb{M}_{\mathcal{U}}$ (Proposition 6.17), there exists some condition $(a,A^{\prime})\leq_{0}(a,A)$ in $\mathbb{M}_{\mathcal{U}}$ which decides the formula $\phi^{*}(\dot{A}_{H},\check{r},\check{\alpha})$, where $\dot{A}_{H}$ is the canonical name for an $\mathbb{M}_{\mathcal{U}}$-generic element of $E^{[\infty]}$. Since $\mathbf{V}[G][\mathcal{U}]\models\mathcal{U}$ is countable and $\mathcal{U}$ is an ultrafilter, there exists some $A_{H}\in[a,A^{\prime}]\cap\mathbf{V}[G]$ such that $A_{H}\leq^{*}B$ for all $B\in\mathcal{U}$. By Proposition 7.9, $A_{H}$ is $\mathbb{M}_{\mathcal{U}}$-generic over $\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$. We shall finish the proof by showing that $a,A_{H}]\subseteq\mathcal{X}$ or $a,A_{H}]\subseteq\mathcal{X}^{c}$.
Suppose that $(a,A^{\prime})\Vdash\phi^{*}(\dot{A}_{H},\check{r},\check{\alpha})$, and let $B\in[a,A_{H}]$. By the Mathias property of $\mathbb{M}_{\mathcal{U}}$, $B$ is also $\mathbb{M}_{\mathcal{U}}$-generic. Since $B\in[a,A^{\prime}]$, the condition $(a,A^{\prime})$ is in the $\mathbb{M}_{\mathcal{U}}$-generic filter generated by $B$, so by the previous claim, $B\in\mathcal{X}$. Therefore, $a,A_{H}]\subseteq\mathcal{X}$. Alternatively, if $(a,A^{\prime})\Vdash\neg\phi^{*}(\dot{A}_{H},\check{r},\check{\alpha})$, then $a,A_{H}]\subseteq\mathcal{X}^{c}$. ∎
We may now prove Corollary 1.8. Given an almost disjoint family $\mathcal{A}\subseteq E^{[\infty]}$, we let:
$\overline{\mathcal{A}}:=\{B\in E^{[\infty]}:B\leq A\text{ for some }A\in\mathcal{A}\}.$
Observe that $\mathcal{H}^{-}(\mathcal{A})\cap\overline{\mathcal{A}}=\emptyset$.
Proof of Corollary 1.8.
Suppose for a contradiction that $\mathcal{A}\in\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$ is a full mad family of block subspaces. Note that $\overline{\mathcal{A}}\in\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$. By Proposition 4.13, $\mathcal{H}(\mathcal{A})$ is a selective coideal, so by Theorem 1.7 we have that $\mathcal{A}$ is $\mathcal{H}$-Ramsey. Let $B\in\mathcal{H}(\mathcal{A})$, and let $C\in\mathcal{H}(\mathcal{A})\mathord{\upharpoonright}A$ be such that $\emptyset,C]\subseteq\overline{\mathcal{A}}$ or $\emptyset,C]\subseteq\overline{\mathcal{A}}^{c}$. Since $C\in\mathcal{H}(\mathcal{A})$ and $\mathcal{H}(\mathcal{A})\cap\overline{\mathcal{A}}=\emptyset$, the first case is not possible. On the other hand, since $\mathcal{A}$ is mad, there exists some $A\in\mathcal{A}$ which is compatible with $C$, i.e. there exists some $D\in E^{[\infty]}$ such that $D\leq C$ and $D\leq A$. Then $D\in[\emptyset,C]\cap\overline{\mathcal{A}}$, another contradiction. ∎
8. Further remarks and open problems
Although Theorem 1.5 asserts the existence of a full mad family of subspaces in $\mathsf{ZFC}$, it remains unknown if every mad family of subspaces in $\mathsf{ZFC}$ is full.
Problem 8.1.
Is there a non-full mad family of subspaces?
While there are various results about the cardinal $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$, the cardinal $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ is not well-understood. In particular, while it is known that $\operatorname{non}(\mathcal{M})\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ holds in $\mathsf{ZFC}$, the $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ variant of the problem remains unsolved.
Problem 8.2.
Is $\operatorname{non}(\mathcal{M})\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ true in $\mathsf{ZFC}$?
Most of the current approaches to the non-existence of mad families of $\omega]^{\omega}$ employ some Ramsey theoretic techniques (such as [M77] [NN18] [ST19]). Consequently, it is difficult to use these approaches to prove the non-existence of mad families of block subspaces over $\mathbb{F}$ for $|\mathbb{F}|>2$. Perhaps the approach highlighted in [T18] may be used to answer the following question:
Problem 8.3.
Let $\kappa$ be an inaccessible cardinal, and let $G$ be $\operatorname{Coll}(\omega,<\!\kappa)$-generic. Are there mad families of block subspaces (over any countable field) in $\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$? What if $\kappa$ is Mahlo?
Acknowledgements
The author would like to express his deepest gratitude to Spencer Unger and Iian Smythe for their helpful comments, suggestions and corrections.